### v2

# On the tails of log-concave density estimators

Abstract.
It is shown that the nonparametric maximum likelihood estimator of a univariate log-concave probability density satisfies some consistency properties in the tail regions.
AMS subject classification:
62G05, 62G07, 62G20, 62G32
Key words:
Chernov bounds, consistency, exponential inequalties, log-linear densities.
1 Introduction
Let $f$ be a log-concave probability density on the real line, that is, $f(x)=\exp(\varphi(x))$ for some concave and upper semicontinuous function $\varphi:\mathbb{R}\to[-\infty,\infty)$. Suppose we observe independent random variables $X_{1},\ldots,X_{n}$ with density $f$ and corresponding distribution function $F$. As noted by Walther (2002) and Pal et al. (2007), for any sample size $n\geq 2$, there exists a unique maximum-likelihood estimator (MLE) $\hat{f}_{n}=\exp(\hat{\varphi}_{n})$ of $f$, where the MLE $\hat{\varphi}_{n}$ of $\varphi$ maximizes
$\sum_{i=1}^{n}\psi(X_{i})$
over all concave functions $\psi:\mathbb{R}\to[-\infty,\infty)$ such that $\int e^{\psi(x)}\,\mathrm{d}x=1$. Denoting the order statistics of $X_{1},\ldots,X_{n}$ with $X_{(1)}<\cdots<X_{(n)}$, this estimator $\hat{\varphi}_{n}$ is piecewise linear on $X_{(1)},X_{(n)}]$ with changes of slope only at observations, and $\hat{\varphi}_{n}=$-$\infty$ outside of $X_{(1)},X_{(n)}]$.
Concerning consistency of $\hat{\varphi}_{n}$, let $\{x\in\mathbb{R}:0<F(x)<1\}=:(a_{o},b_{o})$ with $\infty\leq a_{o}<b_{o}\leq\infty$. It was shown by Dümbgen and Rufibach (2009) that for any fixed interval $a,b]\subset(a_{o},b_{o})$, the supremum of $|\hat{\varphi}_{n}-\varphi|$ over $[a,b]$ is of order $O_{\mathrm{p}}(\rho_{n}^{1/3})$, where $\rho_{n}:=\log(n)/n$. If $\varphi$ is Hölder-continuous on a neighborhood of $[a,b]$ with exponent $\beta\in(1,2]$, this rate improves to $O_{\mathrm{p}}(\rho_{n}^{\beta/(2\beta+1)})$. Uniform consistency of $\hat{\varphi}_{n}$ on arbitrary compact subintervals of $(a_{o},b_{o})$ implies that $\int\bigl{|}\hat{f}_{n}(x)-f(x)\bigr{|}\,\mathrm{d}x\to_{\mathrm{p}}0$. (Throughout this note, asymptotic statements refer to $n\to\infty$.) Pointwise limiting distributions at a single point $x_{o}\in(a_{o},b_{o})$ have been derived by Balabdaoui et al. (2009), assuming that $\varphi$ is twice continuously differentiable in a neighborhood of $x_{o}$. Kim and Samworth (2016) showed that the expected squared Hellinger distance between $\hat{f}_{n}$ and $f$ is of order $O(n^{-4/5})$. Numerous further results about $\hat{\varphi}_{n}$ and $\hat{f}_{n}$ have been derived thereafter, including multivariate settings, see the review of Samworth (2018). In the present univariate setting, fast algorithms for the computation of $\hat{\varphi}_{n}$ and related objects are provided by Dümbgen and Rufibach (2011) and Dümbgen et al. (2021). Experiments with simulated data show that even in the tail regions, that is, close to $a_{o}$ and $b_{o}$, the estimator $\hat{\varphi}_{n}$ is surprisingly accurate. In view of this empirical finding, Müller and Rufibach (2009) developed new estimators for extreme value analysis with excellent empirical performance. However, the currently available theory about the asymptotic properties of $\hat{\varphi}_{n}$ does not explain its good performance in the tail regions. The present note provides a first step into that direction by proving several results about $\hat{\varphi}_{n}(x)$ and the right-sided derivative $\hat{\varphi}_{n}^{\prime}(x+)$ for $x$ close to $b_{o}$. By symmetry, these findings carry over to results in the left tail region. The main results are presented in Section 2 while Sections 3 and 4 provide the proofs.
2 Main results
We start with a simple consequence of the pointwise consistency of $\hat{f}_{n}$, $\hat{\varphi}_{n}$ and concavity of $\varphi,\hat{\varphi}_{n}$.
Theorem 1.
The estimator $\hat{f}_{n}$ does not overestimate $f$ in the sense that
$\sup_{x\in\mathbb{R}}\,\bigl{(}\hat{f}_{n}(x)-f(x)\bigr{)}^{+}\ \to_{\mathrm{p}}\ 0.$
Moreover, for any sequence $(b_{n})_{n}$ in $(a_{o},b_{o})$ with limit $b_{o}$,
$\hat{\varphi}_{n}^{\prime}(b_{n}+)\ \begin{cases}\leq\ \varphi^{\prime}(b_{o}-)+o_{\mathrm{p}}(1)&\text{if}\ \varphi^{\prime}(b_{o}-)>-\infty,\\ \to_{\mathrm{p}}\ -\infty&\text{if}\ \varphi^{\prime}(b_{o}-)=-\infty,\end{cases}$
where $\hat{\varphi}_{n}^{\prime}(x+):=$-$\infty$ for $x\geq X_{(n)}$.
The remaining goal is to show that the right tails are not “severely underestimated”, and for this task we distinguish the cases $b_{o}=\infty$ and $b_{o}<\infty$.
Theorem 2.
Suppose that $b_{o}<\infty$.
(a)  Let $f(b_{o})=0$. Then for any fixed $a\in(a_{o},b_{o})$,
$\sup_{x\geq a}\bigl{|}\hat{f}_{n}(x)-f(x)\bigr{|}\ \to_{\mathrm{p}}\ 0.$
Moreover, for any given sequence $(b_{n})_{n}$ in $(a_{o},b_{o})$ with limit $b_{o}$,
$\hat{\varphi}_{n}^{\prime}(b_{n}+)\ \to_{\mathrm{p}}\ \varphi^{\prime}(b_{o}-)=-\infty.$
(b)  Let $f(b_{o})>0$. Then for arbitrary fixed intervals $a,b_{n}]\subset(a_{o},b_{o})$ such that $b_{n}\uparrow b_{o}$ and $n(1-F(b_{n}))\to\infty$,
$\sup_{x\in[a,b_{n}]}\bigl{|}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{|}\ \to_{\mathrm{p}}\ 0.$
(c)  Let $f(b_{o})>0$ and $\varphi^{\prime}(b_{o}-)>$-$\infty$. Then for any given sequence $(b_{n})_{n}$ in $(a_{o},b_{o})$ such that $b_{n}\uparrow b_{o}$ and $\rho_{n}^{-1/3}(1-F(b_{n}))\to\infty$,
$\hat{\varphi}_{n}^{\prime}(b_{n}+)\ \to_{\mathrm{p}}\ \varphi^{\prime}(b_{o}-).$
Example 1.
We illustrate Theorem 1 and Theorem 2 (b-c) for samples from the uniform distribution on $[0,1]$. Figure 1 depicts the functions $\hat{\varphi}_{n}$ (left panel) and $\hat{\varphi}_{n}^{\prime}(\cdot+)$ (right panel) for one “typical sample” of size $n=150$ (top), $n=500$ (middle) and $n=2000$ (bottom). Figure 2 shows the performance of $\hat{\varphi}_{n}$ in $10000$ simulations of a sample of size $n=150$ (top), $n=500$ (middle) and $n=2000$ (bottom). The left panels show the estimated $\gamma$-quantiles of $\hat{\varphi}_{n}(x)$, $x\in(0,1)$, for $\gamma=0.01,0.1,0.25,0.5,0.75,0.9,0.99$. For the same values of $\gamma$, the right panels show the estimated $\gamma$-quantiles of $\hat{\varphi}_{n}^{\prime}(x+)$, $x\in(0,1)$. As expected, the estimators $\hat{\varphi}_{n}$ and $\hat{\varphi}_{n}^{\prime}(\cdot+)$ suffer from a substantial bias very close to the boundaries 0 and $1$, but these problematic regions shrink as the sample size $n$ increases.
Figure 1: The functions $\hat{\varphi}_{n}$ (left panel) and $\hat{\varphi}_{n}^{\prime}(\cdot+)$ (right panel) for one particular sample of size $n=150$ (top), $n=500$ (middle) and $n=2000$ (bottom) from $\mathrm{Unif}[0,1]$. The sample is indicated as a rug plot, and the true values $\varphi$ and $\varphi^{\prime}$ are shown in red.
Figure 2: Estimated $\gamma$-quantiles of $\hat{\varphi}_{n}(x)$ (left panel) and $\hat{\varphi}_{n}^{\prime}(x+)$ (right panel) for samples of size $n=150$ (top), $n=500$ (middle) and $n=2000$ (bottom) from $\mathrm{Unif}[0,1]$. The true values $\varphi(x)$ and $\varphi^{\prime}(x)$ are shown in red.
If the support of $P$ is unbounded to the right, then $\varphi(x)\to$-$\infty$ and $\varphi^{\prime}(x+)\to\varphi^{\prime}(\infty-)\in[-\infty,0)$ as $x\to\infty$. Here are some results complementing Theorem 1.
Theorem 3.
Suppose that $b_{o}=\infty$. Let $(b_{n})_{n}$ be a sequence in $(a_{o},\infty)$ such that $b_{n}\to\infty$ and $(1-F(b_{n}))/\rho_{n}\to\infty$.
(a)  With asymptotic probability one, $\hat{f}_{n}(b_{n})>0$, and
$\hat{\varphi}_{n}^{\prime}(b_{n}+)\ \to_{\mathrm{p}}\ \varphi^{\prime}(\infty-).$
(b)  Suppose that $\varphi^{\prime}(\infty-)>$-$\infty$. Then for any fixed $a\in(a_{o},\infty)$,
$\max_{x\in[a,b_{n}]}\,\frac{\bigl{|}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{|}}{1+|\varphi(x)|}\ \to_{\mathrm{p}}\ 0.$
(c)  Suppose that $\varphi^{\prime}(\infty-)=$-$\infty$ and $\varphi$ is differentiable on some halfline $(a_{*},\infty)\subset(a_{o},\infty)$ with Lipschitz-continuous derivative $\varphi^{\prime}$. Then, for arbitary fixed $a\in(a_{*},\infty)$ such that $\varphi^{\prime}(a)<0$,
$\sup_{x\in[a,b_{n}]}\,\Bigl{(}\frac{\hat{\varphi}_{n}^{\prime}(x+)}{\varphi^{\prime}(x)}-1\Bigl{)}^{+}\$$\to_{\mathrm{p}}\ 0,$
$\sup_{x\in[a,b_{n}]}\,\frac{\bigl{(}\varphi(x)-\hat{\varphi}_{n}(x)\bigr{)}^{+}}{1+|\varphi(x)|}\$$\to_{\mathrm{p}}\ 0,$
$\sup_{x\in[a,b_{n}]}\,\bigl{(}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{)}^{+}\$$\to_{\mathrm{p}}\ 0.$
Parts (a) and (b) apply, for instance, if $P$ is a logistic distribution or a gamma distribution with shape parameter in $1,\infty)$. Parts (a) and (c) explain why the estimator $\hat{\varphi}_{n}$ is remarkably accurate in the tails if, for instance, $P$ is a Gaussian distribution. In particular, for any fixed $\varepsilon>0$,
$\operatorname{\mathbb{P}}\bigl{(}\hat{f}_{n}(x)\leq(1+\varepsilon)f(x)\ \text{for all}\ x\in[a,b_{n}]\bigr{)}\ \to\ 1,$
which is substantially stronger than the first conclusion of Theorem 1. Furthermore, since $\varphi^{\prime}<0$ on $a,b_{n}]$,
$\operatorname{\mathbb{P}}\bigl{(}\hat{\varphi}_{n}^{\prime}(x+)\geq(1+\varepsilon)\varphi^{\prime}(x)\ \text{for all}\ x\in[a,b_{n}]\bigr{)}\ \to\ 1,$
which complements the second conclusion of Theorem 1.
Example 2.
We illustrate Theorem 1 and Theorem 3 (b) for samples from the standard Gaussian distribution. Figure 3 depicts the functions $\hat{\varphi}_{n}$ (left panel) and $\hat{\varphi}_{n}^{\prime}(\cdot+)$ (right panel) for one “typical sample” of size $n=150$ (top), $n=500$ (middle) and $n=2000$ (bottom). Figure 4 shows the performance of $\hat{\varphi}_{n}$ in $10000$ simulations of a sample of size $n=150$ (top), $n=500$ (middle) and $n=2000$ (bottom). The left-hand side shows the estimated $\gamma$-quantiles of $\hat{\varphi}_{n}(x)$, $x\in(-4,4)$ for $\gamma=0.01,0.1,0.25,0.5,0.75,0.9,0.99$. For the same values of $\gamma$, the right-hand side shows the estimated $\gamma$-quantiles of $\hat{\varphi}_{n}^{\prime}(x+)$, $x\in(-4,4)$.
Figure 3: The functions $\hat{\varphi}_{n}$ (left panel) and $\hat{\varphi}_{n}^{\prime}(\cdot+)$ (right panel) for one particular sample of size $n=150$ (top), $n=500$ (middle) and $n=2000$ (bottom) from $\mathrm{N}(0,1)$. The sample is indicated as a rug plot, and the true values $\varphi$ and $\varphi^{\prime}$ are shown in red.
Figure 4: Estimated $\gamma$-quantiles of $\hat{\varphi}_{n}(x)$ (left panel) and $\hat{\varphi}_{n}^{\prime}(x+)$ (right panel) for samples of size $n=150$ (top), $n=500$ (middle) and $n=2000$ (bottom) from $\mathrm{N}(0,1)$. The true values $\varphi(x)$ and $\varphi^{\prime}(x)$ are shown in red.
3 Auxiliary results
In what follows, let $P$ and $\hat{P}_{n}$ be the distribution with density $f$ and $\hat{f}_{n}$, respectively. The corresponding distribution functions are denoted by $F$ and $\hat{F}_{n}$, respectively. In addition, let $\hat{P}_{n}^{\mathrm{emp}}$ and $\hat{F}_{n}^{\mathrm{emp}}$ be the empirical distribution and the empirical distribution function, respectively, of the observations $X_{1},\ldots,X_{n}$.
3.1 More about $\hat{f}_{n}$ and $\hat{\varphi}_{n}$
We mentioned already some properties of $\hat{\varphi}_{n}$ and $\hat{f}_{n}$. Recall that for any fixed $a,b]\subset(a_{o},b_{o})$,
$\sup_{x\in[a,b]}\,\bigl{|}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{|}\ \to_{\mathrm{p}}\ 0,$ (1)
and since $\varphi$ is bounded on $[a,b]$, this implies that
$\sup_{x\in[a,b]}\,\bigl{|}\hat{f}_{n}(x)-f(x)\bigr{|}\ \to_{\mathrm{p}}\ 0.$ (2)
An important consequence of the latter result is that
$\sup_{B\in\mathrm{Borel}(\mathbb{R})}\bigl{|}\hat{P}_{n}(B)-P(B)\bigr{|}\ =\ 2^{-1}\int\bigl{|}\hat{f}_{n}(x)-f(x)\bigr{|}\,\mathrm{d}x\ \to_{\mathrm{p}}\ 0,$ (3)
see Dümbgen and Rufibach (2009). The latter paper also provides the following key inequalities: Let
$\hat{\mathcal{S}}_{n}\ :=\ \{X_{(1)},X_{(n)}\}\cup\bigl{\{}x\in(X_{(1)},X_{(n)}):\hat{\varphi}_{n}(x-)>\hat{\varphi}_{n}(x+)\bigr{\}},$
the set of kinks of $\hat{\varphi}_{n}$. Then for arbitrary $b\in\mathbb{R}$,
$\int(x-b)^{+}\,\hat{P}_{n}^{\mathrm{emp}}(\mathrm{d}x)\ \begin{cases}\displaystyle\geq\ \int(x-b)^{+}\,\hat{P}_{n}(\mathrm{d}x),\\ \displaystyle=\ \int(x-b)^{+}\,\hat{P}_{n}(\mathrm{d}x)&\text{if}\ b\in\hat{\mathcal{S}}_{n}.\end{cases}$ (4)
Moreover, for $b\in\hat{\mathcal{S}}_{n}$,
$\hat{F}_{n}^{\mathrm{emp}}(b-)\ \leq\ \hat{F}_{n}(b)\ \leq\ \hat{F}_{n}^{\mathrm{emp}}(b)$ (5)
Finally,
$\int x\,\hat{P}_{n}^{\mathrm{emp}}(\mathrm{d}x)\ =\ \int x\,\hat{P}_{n}(\mathrm{d}x).$
3.2 Inequalities for $\hat{P}_{n}^{\mathrm{emp}}$
Concerning $\hat{F}_{n}^{\mathrm{emp}}$, note the following useful inequality: For any $b<b_{o}$,
$\operatorname{\mathbb{E}}\biggl{(}\sup_{x\leq b}\,\Bigl{|}\frac{\hat{F}_{n}^{\mathrm{emp}}(x)-F(x)}{1-F(x)}\Bigr{|}^{2}\biggr{)}\ \leq\ \frac{4}{n(1-F(b))}.$
This follows from the well-known fact that $M_{x}:=[\hat{F}_{n}^{\mathrm{emp}}(x)-F(x)]/[1-F(x)]$ defines a martingale $(M_{x})_{x<b_{o}}$ and from one of Doob’s martingale inequalities (Shorack and Wellner, 1986; Hall and Heyde, 1980). In particular, for any sequence of numbers $b_{n}\in(a_{o},b_{o})$,
$\sup_{x\leq b_{n}}\Bigl{|}\frac{1-\hat{F}_{n}^{\mathrm{emp}}(x)}{1-F(x)}-1\Bigr{|}\ \to_{\mathrm{p}}\ 0\quad\text{if}\ n(1-F(b_{n}))\to\infty.$ (6)
Combining this with (5) leads to the fact that
$\max_{x\in\hat{\mathcal{S}}_{n}:x\leq b_{n}}\Bigl{|}\frac{1-\hat{F}_{n}(x)}{1-F(x)}-1\Bigr{|}\ \to_{\mathrm{p}}\ 0\quad\text{if}\ n(1-F(b_{n}))\to\infty.$ (7)
Here is another useful result about $\hat{P}_{n}^{\mathrm{emp}}(I)-P(I)$ over real intervals $I$.
Proposition 1 (Consistency of $\hat{P}_{n}^{\mathrm{emp}}$).
For any constant $\tau>2$, with asymptotic probability one,
$\bigl{|}\hat{P}_{n}^{\mathrm{emp}}(I)-P(I)\bigr{|}\ \leq\ \sqrt{2\tau\rho_{n}\min\{\hat{P}_{n}^{\mathrm{emp}}(I),P(I)\}}+(\tau+2)\rho_{n}$
for arbitrary intervals $I\subset\mathbb{R}$.
Proof of Proposition 1.
Note first that for arbitrary indices $0\leq j<k\leq n+1$ with $k-j\leq n$, the random variable $P(X_{(j)},X_{(k)})$ follows a beta distribution with parameters kjk-j and n+1k+jn+1-k+j, see Chapter 3.1 of Shorack and Wellner (1986). In particular, its mean equals
$p_{njk}\ :=\ \frac{k-j}{n+1},$
and it follows from Proposition 2.1 of Dümbgen (1998) that for any $\eta>0$,
$\operatorname{\mathbb{P}}\bigl{[}\Psi\bigl{(}P(X_{(j)},X_{(k)}),p_{njk}\bigr{)}\geq\eta\rho_{n}\bigr{]}\ \leq\ 2\exp(-(n+1)\eta\rho_{n})\ <\ 2n^{-\eta},$
where $\Psi(x,p):=p\log(p/x)+(1-p)\log[(1-p)/(1-x)]$. Moreover, for $c>0$, the inequality $\Psi(x,p)<c$ implies that $|x-p|<\sqrt{2cp(1-p)}+|1-2p|c$. Furthermore, since $\Psi(x,p)=K_{0}(x,p)=K_{1}(p,x)$ in the notation of Dümbgen and Wellner (2023), Lemma S.12 in the latter paper shows that $|x-p|<\sqrt{2cx}+(2/3)|1-2x|c$ whenever $\Psi(x,p)<c$. Consequently, the probability that
$\bigl{|}P(X_{(j)},X_{(k)})-p_{njk}\bigr{|}\ \leq\ \sqrt{2\eta\rho_{n}\min\{P(X_{(j)},X_{(k)}),p_{njk}\}}+\eta\rho_{n}$ (8)
for all indices $0\leq j<k\leq n+1$ is at least $1-2\binom{n+2}{2}n^{-\eta}=1-(1+3/n+2/n^{2})n^{2-\eta}$, and this converges to $1$ if $\eta>2$.
It remains to be shown that in case of $\eta\in(2,\tau)$, inequality (8) implies that with asymptotic probability one,
$\bigl{|}P(I)-\hat{P}_{n}^{\mathrm{emp}}(I)\bigr{|}\ \leq\ \sqrt{2\tau\rho_{n}\min\{P(I),\hat{P}_{n}^{\mathrm{emp}}(I)\}}+(\tau+2)\rho_{n}$
for all intervals $I\subset\mathbb{R}$. Note first that
$\max_{\ell=1,\ldots,n+1}P(X_{(\ell-1)},X_{(\ell)})\ \leq\ \rho_{n}+O_{\mathrm{p}}(n^{-1}).$ (9)
This can be deduced from the well-known representation $P(X_{(\ell-1)},X_{(\ell)})=E_{\ell}/\sum_{i=1}^{n+1}E_{i}$ with independent, standard exponential random variables $E_{1},\ldots,E_{n+1}$. Now, for an arbitrary nonvoid interval $I$, let $j=j_{n}(I)\in\{0,1,\ldots,n\}$ be maximal and $k=k_{n}(I)\in\{1,2,\ldots,n+1\}$ be minimal such that $I\subset[X_{(j)},X_{(k)}]$. If $k-j\leq 2$, then
$\hat{P}_{n}^{\mathrm{emp}}(I)-P(I)\ \leq\ \hat{P}_{n}^{\mathrm{emp}}(I)\ \leq\ 3/n\ =\ o(\rho_{n})$
and
[…]
In case of $k-j\geq 3$, let $\tilde{I}:=[X_{(j+1)},X_{(k-1)}]\subset I$ with
$\hat{P}_{n}^{\mathrm{emp}}(\tilde{I})=\frac{k-j-1}{n}\ >\ \frac{k-j-2}{n+1}=p_{n,j+1,k-1}.$
Consequently, it follows from (8) and (9) that
$P(I)-\hat{P}_{n}^{\mathrm{emp}}(I)\$$\leq\ P(\tilde{I})-\hat{P}_{n}^{\mathrm{emp}}(\tilde{I})+2\max_{\ell=0,\ldots,n}P(X_{(\ell-1)},X_{(\ell)})$
$\leq\ P(\tilde{I})-p_{n,j+1,k-1}+(2+o_{\mathrm{p}}(1))\rho_{n}$
$\leq\ \sqrt{2\eta\rho_{n}\min\{P(\tilde{I}),p_{n,j+1,k-1}\}}+(\eta+2+o_{\mathrm{p}}(1))\rho_{n}$
$\leq\ \sqrt{2\eta\rho_{n}\min\{P(I),\hat{P}_{n}^{\mathrm{emp}}(I)\}}+(\eta+2+o_{\mathrm{p}}(1))\rho_{n},$
and
[…]
$\leq\ o(\rho_{n})+(1+1/n)\Bigl{(}\sqrt{2\eta\rho_{n}\min\{P(I),\hat{P}_{n}^{\mathrm{emp}}(I)\}}+\eta\rho_{n}\Bigr{)}$
$=\ \sqrt{2(\eta+o(1))\rho_{n}\min\{P(I),\hat{P}_{n}^{\mathrm{emp}}(I)\}}+(\eta+o(1))\rho_{n},$
where the terms $o_{\mathrm{p}}(1)$ and $o(1)$ depend only on $n$, not on the interval $I$. ∎
3.3 Truncated and conditional means of $P$
For $\infty<a<b\leq\infty$, let
$M(a,b)\$$:=\ \int_{(a,b)}(x-a)\,P(\mathrm{d}x),$
$\mu(a,b)\$$:=\ \begin{cases}P(a,b)^{-1}M(a,b)&\text{if}\ P(a,b)>0,\\ 0&\text{else}.\end{cases}$
For $\infty\leq a<b<\infty$, we set
[…]
Further, let
$\mu(a)\ :=\ \mu(a,\infty).$ (10)
The univariate function $\mu$ is known as the mean excess function or mean residual lifetime in fields such as extreme value theory and actuarial science.
To formulate various approximations and inequalities for these functions $M$, $\mu$ and $W$, we need three auxiliary functions and some properties thereof.
Proposition 2.
Let $N,\nu,V:\mathbb{R}\to(0,\infty)$ be given by
[…]
These functions are continuously differentiable with $N^{\prime},\nu^{\prime},V^{\prime}>0$, where $N(0)=\nu(0)=V(0)=1/2$, $N^{\prime}(0)=1/3$, $\nu^{\prime}(0)=1/12$ and $V^{\prime}(0)=1/6$. Moreover,
$\lim_{t\to-\infty}\,t^{2}N(t)\ =\ \lim_{t\to-\infty}\,|t|\nu(t)\ =\ \lim_{t\to-\infty}\,|t|V(t)\ =\ 1.$
This proposition follows from elementary calculus. The limits of $t^{2}N(t)$, $|t|\nu(t)$ and $|t|V(t)$ as $t\to$-$\infty$ follow from the explicit formulae
[…]
for $t\neq 0$.
The next proposition summarizes several useful properties of the functions $M$, $\mu$ and $W$ and their relation to $f$ and $\varphi$. The monotonicity property of $\mu$ in part (a) was noted already by Bagnoli and Bergstrom (2005). In the proposition’s proof and later on, we use repeatedly well-known results about the stochastic and likelihood ratio orders between probability distributions on the real line, see Shaked and Shanthikumar (2007) for the foundations. Specifically, let $P_{1}$ and $P_{2}$ be probability distributions on the real line with densities $f_{1}$ and $f_{2}$, respectively. If $f_{1}\geq f_{2}$ on $(-\infty,x_{o})$ and $f_{1}\leq f_{2}$ on $(x_{o},\infty)$ for some real number $x_{o}$, then $P_{1}\leq_{\rm st}P_{2}$, where $\leq_{\rm st}$ denotes stochastic order. In particular, if $f_{2}/f_{1}$ is non-decreasing on $\{f_{1}+f_{2}>0\}$, then such a number $x_{o}$ has to exist, whence $P_{1}\leq_{\rm st}P_{2}$.
Proposition 3 (Properties of $M$ and $\mu$).
(a)  The function $\mu$ given by (10) is non-increasing and Lipschitz-continuous with constant one.
(b)  Let $\infty<a<b<c\leq\infty$ such that $P(a,b)>0$. Then,
$\frac{1}{p}\ \leq\ \frac{M(a,c)}{M(a,b)}\ \leq\ \frac{1}{p+(1-p)\log(1-p)},$
where $p:=P(a,b)/P(a,c)\in(0,1]$ and $0\log 0:=0$.
(c)  For arbitrary real numbers $a<b$ in $a_{o},b_{o}]$,
$f(a)(b-a)^{2}N\bigl{(}\varphi^{\prime}(b-)(b-a)\bigr{)}\ \leq\ M(a,b)\ \leq\ f(a)(b-a)^{2}N\bigl{(}\varphi^{\prime}(a+)(b-a)\bigr{)},$
$f(a)(b-a)^{2}V\bigl{(}\varphi^{\prime}(b-)(b-a)\bigr{)}\ \leq\ W(a,b)\ \leq\ f(a)(b-a)^{2}V\bigl{(}\varphi^{\prime}(a+)(b-a)\bigr{)},$
where $N(-\infty),V(-\infty):=0$ and $N(\infty),V(\infty):=\infty$. Moreover,
$(b-a)\nu\bigl{(}\varphi^{\prime}(b-)(b-a)\bigr{)}\ \leq\ \mu(a,b)\ \leq\ (b-a)\nu\bigl{(}\varphi^{\prime}(a+)(b-a)\bigr{)}$
where $\nu(-\infty):=0$, $\nu(\infty):=1$.
(d)  Suppose that $b_{o}=\infty$. Then for arbitrary real $a\in[a_{o},\infty)$ with $\varphi^{\prime}(a+)<0$,
$\frac{f(a)}{\varphi^{\prime}(\infty-)^{2}}\ \leq\ M(a,\infty)\ \leq\ \frac{f(a)}{\varphi^{\prime}(a+)^{2}}\quad\text{and}\quad\frac{1}{|\varphi^{\prime}(\infty-)|}\ \leq\ \mu(a)\ \leq\ \frac{1}{|\varphi^{\prime}(a+)|}.$
(e)  Suppose that $\varphi$ is differentiable on some interval $(a,b)\subset[a_{o},b_{o}]$ with $a\in\mathbb{R}$ and $\varphi^{\prime}(a+)<0$. Further, let $\varphi^{\prime}$ be Lipschitz-continuous with constant $L$ on $(a,b)$. Then
[…]
and
$\mu(a,b)\ \geq\ \exp\bigl{(}-3L\varphi^{\prime}(a+)^{-2}\bigr{)}\cdot\begin{cases}\displaystyle\frac{1}{|\varphi^{\prime}(a+)|}&\text{if}\ b=\infty,\\[8.61108pt] (b-a)\nu\bigl{(}\varphi^{\prime}(a+)(b-a)\bigr{)}&\text{if}\ b<\infty.\end{cases}$
Proof of Proposition 3.
As to part (a), by means of Fubini’s theorem we may write
$\mu(a)\ =\ (1-F(a))^{-1}\int\int_{0}^{\infty}1_{[r<y-a]}\,\mathrm{d}r\,P(\mathrm{d}y)\ =\ \int_{0}^{\infty}\frac{1-F(a+r)}{1-F(a)}\,\mathrm{d}r$
for $a\in\mathbb{R}$ with $F(a)<1$. For $a<a^{\prime}$ with $F(a^{\prime})=1$,
$\mu(a)-\mu(a^{\prime})\ =\ \mu(a)\ =\ \int_{0}^{a^{\prime}-a}\frac{1-F(a+r)}{1-F(a)}\,\mathrm{d}r\ \in\ [0,a^{\prime}-a].$
For $a<a^{\prime}$ with $F(a^{\prime})<1$,
$\mu(a)-\mu(a^{\prime})\$$=\ \int_{0}^{\infty}\frac{1-F(a+r)}{1-F(a)}\,\mathrm{d}r-\int_{0}^{\infty}\frac{1-F(a^{\prime}+r)}{1-F(a^{\prime})}\,\mathrm{d}r$
$\leq\ \int_{0}^{\infty}\frac{1-F(a+r)}{1-F(a)}\,\mathrm{d}r-\int_{0}^{\infty}\frac{1-F(a^{\prime}+r)}{1-F(a)}\,\mathrm{d}r$
$=\ \int_{a}^{a^{\prime}}\frac{1-F(a+r)}{1-F(a)}\,\mathrm{d}r\ \leq\ a^{\prime}-a.$
That $\mu(a)\geq\mu(a^{\prime})$ was noted already by Bagnoli and Bergstrom (2005), but for the reader’s convenience, we provide an argument here. For $\xi\in\{a,a^{\prime}\}$, one may write
$\mu(\xi)\ =\ \int_{[0,\infty)}zf_{\xi}(z)\,\mathrm{d}z$
with the probability density $f_{\xi}$ on $0,\infty)$ given by $f_{\xi}(z)\ :=\ \exp\bigl{(}\varphi(\xi+z)-\log(1-F(\xi))\bigr{)}$. By concavity of $\varphi$, $f_{a}/f_{a^{\prime}}$ is non-decreasing on $\{f_{a}>0\}=[0,b_{o}-a)\supset\{f_{a^{\prime}}>0\}$. This implies that the distribution with density $f_{a}$ is stochastically greater than (or equal to) the distribution with density $f_{a^{\prime}}$. In particular, the mean $\mu(a)$ of the former is not smaller than the mean $\mu(a^{\prime})$ of the latter.
To prove part (b), it suffices to consider the nontrivial case that $P(b,c)>0$, i.e. $p\in(0,1)$. Note that the ratios $M(a,c)/M(a,b)$, $\mu(a,c)/\mu(a,b)$ and $p=P(a,b)/P(a,c)$ remain the same if we replace $P$ with the conditional distribution $P(a,c)^{-1}P(\cdot\cap(a,c))$. Thus we may assume that $P(a,c)=1$, and we may replace $c$ with $\infty$. With a random variable $X$ with this (modified) distribution $P$, we may write
$\frac{M(a,c)}{M(a,b)}\ =\ \frac{\operatorname{\mathbb{E}}(X-a)}{\operatorname{\mathbb{E}}(1_{[X\leq b]}(X-a))},\quad\frac{\mu(a,c)}{\mu(a,b)}\ =\ \frac{\operatorname{\mathbb{E}}(X-a)}{\operatorname{\mathbb{E}}(X-a\,|\,X\leq b)}\ =\ \frac{pM(a,c)}{M(a,b)}.$
Now let $\tilde{f}(x):=1_{[x>a]}\lambda\exp(-\lambda(x-a))$ with $\lambda>0$ such that $\int_{a}^{b}\tilde{f}(x)\,\mathrm{d}x=p$, that is,
$\lambda\ =\ \frac{-\log(1-p)}{b-a}.$
By concavity of $\log f$ and linearity of $\log\tilde{f}$ on $(a,\infty)$, either $f\equiv\tilde{f}$, or there exist numbers $a\leq x_{1}<x_{2}\leq\infty$ such that
[…]
Note that $x_{1}=a$ would imply that $f>\tilde{f}$ on $(a,b)$ or $f<\tilde{f}$ on $(b,\infty)$, and in both cases we would end up with $\int_{a}^{b}f(x)\,\mathrm{d}x\neq\int_{a}^{b}\tilde{f}(x)\,\mathrm{d}x$. Similiarly one can exclude the cases $x_{2}=\infty$, $x_{1}\geq b$ and $x_{2}\leq b$. Consequently, we know that there exist constants $a<x_{1}<b<x_{2}<\infty$ such that
[…]
In particular, if $\tilde{X}$ is a random variable with density $\tilde{f}$, then
$\mathcal{L}(X\,|\,X\leq b)\ \geq_{\rm st}\ \mathcal{L}(\tilde{X}\,|\,\tilde{X}\leq b)\quad\text{and}\quad\mathcal{L}(X\,|\,X>b)\ \leq_{\rm st}\ \mathcal{L}(\tilde{X}\,|\,\tilde{X}>b),$
where $\mathcal{L}(\cdot)$ stands for ‘distribution’. In particular,
$\frac{M(a,c)}{M(a,b)}\$= $=\ 1+\frac{(1-p)\operatorname{\mathbb{E}}(X-a\,|\,X>b)}{p\operatorname{\mathbb{E}}(X-a\,|\,X\leq b)}$
$\leq\ 1+\frac{(1-p)\operatorname{\mathbb{E}}(\tilde{X}-a\,|\,\tilde{X}>b)}{p\operatorname{\mathbb{E}}(\tilde{X}-a\,|\,\tilde{X}\leq b)}\ =\ \frac{\operatorname{\mathbb{E}}(\tilde{X}-a)}{\operatorname{\mathbb{E}}(1_{[\tilde{X}\leq b]}(\tilde{X}-a))}.$
Elementary calculations reveal that the latter ratio is equal to $1/[p+(1-p)\log(1-p)]$, which yields our upper bound for $M(a,c)/M(a,b)$. Concerning the lower bound, note that
$\frac{M(a,c)}{M(a,b)}\ =\ 1+\frac{(1-p)\operatorname{\mathbb{E}}(X-a\,|\,X>b)}{p\operatorname{\mathbb{E}}(X-a\,|\,X\leq b)}\ \geq\ 1+\frac{(1-p)}{p}\ =\ \frac{1}{p}.$
To prove part (c), we consider up to three functions $\psi_{1},\psi_{2},\psi_{3}:[0,b-a]\to[-\infty,\infty]$ given by
$\psi_{1}(z)\$$:=\ \varphi^{\prime}(b-)z\qquad(\text{if}\ \varphi^{\prime}(b_{o}-)>-\infty),$
$\psi_{2}(z)\$$:=\ \varphi(a+z)-\varphi(a),$
$\psi_{3}(z)\$$:=\ \varphi^{\prime}(a+)z\qquad(\text{if}\ \varphi^{\prime}(a+)<\infty).$
Concavity of $\varphi$ implies that $\psi_{1}\leq\psi_{2}\leq\psi_{3}$, so
$\frac{M(a,b)}{f(a)}\ =\ \int_{(a,b)}(x-a)\exp(\psi_{2}(x-a))\,\mathrm{d}x\ =\ \int_{0}^{b-a}z\exp(\psi_{2}(z))\,\mathrm{d}z$
lies between
$\int_{0}^{b-a}z\exp(\psi_{1}(z))\,\mathrm{d}z\ =\ (b-a)^{2}N\bigl{(}\varphi^{\prime}(b-)(b-a)\bigr{)}$
and
$\int_{0}^{b-a}z\exp(\psi_{3}(z))\,\mathrm{d}z\ =\ (b-a)^{2}N\bigl{(}\varphi^{\prime}(a+)(b-a)\bigr{)}.$
Analogously, one can show that $W(a,b)/f(a)$ lies between $(b-a)^{2}V\bigl{(}\varphi^{\prime}(b-)(b-a)\bigr{)}$ and $(b-a)^{2}V\bigl{(}\varphi^{\prime}(a+)(b-a)\bigr{)}$.
Concerning the inequalities for $\mu(a,b)$, note that in case of $\varphi^{\prime}(b-)>$-$\infty$,
$\psi_{2}(z)-\psi_{1}(z)\ =\ \int_{0}^{z}\bigl{[}\varphi^{\prime}((a+t)+)-\varphi^{\prime}(b-)\bigr{]}\,\mathrm{d}t$
is non-decreasing in $z\in[0,b-a]$, so the probability distribution $P_{1}$ on [0,ba][0,b-a] with density proportional to $\exp(\psi_{1})$ is stochastically smaller than (or equal to) the probability distribution $P_{2}$ on [0,ba][0,b-a] with density proportional to $\exp(\psi_{2})$. Thus,
$\mu(a,b)\ =\ \int_{[0,b-a]}z\,P_{2}(\mathrm{d}z)\ \geq\ \int_{[0,b-a]}z\,P_{1}(\mathrm{d}z)\ =\ (b-a)\nu\bigl{(}\varphi^{\prime}(b-)(b-a)\bigr{)}.$
Analogously, if $\varphi^{\prime}(a+)<\infty$, then $\psi_{3}-\psi_{2}$ is non-decreasing on [0,ba][0,b-a], and this implies that
$\mu(a,b)\ \leq\ (b-a)\nu\bigl{(}\varphi^{\prime}(a+)(b-a)\bigr{)}.$
Part (d) is verified similarly as part (c). Here we consider two or three probability densities $f_{1},f_{2},f_{3}$ on $0,\infty)$ given by $f_{1}(z):=\lambda_{1}\exp(-\lambda_{1}z)$ with $\lambda_{1}:=-\varphi^{\prime}(\infty-)$ (if $\varphi^{\prime}(\infty-)>$-$\infty$), $f_{2}(z):=\exp\bigl{(}\varphi(a+z)-\log(1-F(a))\bigr{)}$ and $f_{3}(z):=\lambda_{3}\exp(-\lambda_{3}z)$ with $\lambda_{3}:=-\varphi^{\prime}(a+)$. If $\varphi^{\prime}(\infty-)>$-$\infty$, then $f_{2}/f_{1}$ is non-decreasing, whence $\mu(a)\geq 1/\lambda_{1}=1/|\varphi^{\prime}(\infty-)|$. And $f_{3}/f_{2}$ is non-decreasing too, so $\mu(x)\leq 1/\lambda_{3}=1/|\varphi^{\prime}(a+)|$.
As to part (e), it suffices to prove the inequalities for $M(a,b)$ and $\mu(a,b)$ with $b<\infty$, because $t:=\varphi^{\prime}(a+)(b-a)\to$-$\infty$ and $t^{2}N(t)\to 1$, $|t|\nu(t)\to 1$ as $b\to\infty$, see Proposition 2. For $z\in[0,b-a]$, let $\psi_{2}(z)=\varphi(a+z)-\varphi(a)$ as before and
$\psi_{4}(z)\ :=\ \varphi^{\prime}(a+)z-Lz^{2}/2.$
The difference $\psi_{2}-\psi_{4}$ is non-negative and non-decreasing, because
$\psi_{2}(z)-\psi_{4}(z)\ =\ \int_{0}^{z}\bigl{[}\varphi^{\prime}(a+t)-\varphi^{\prime}(a+)+Lt\bigr{]}\,\mathrm{d}t,$
and the integrand is non-negative by Lipschitz-continuity of $\varphi^{\prime}$ on $(a,b)$ with constant $L$. Hence, $M(a,b)$ is not smaller than
$f(a)\int_{0}^{b-a}z\exp(\psi_{4}(z))\,\mathrm{d}z\$$=\ f(a)\int_{0}^{b-a}z\exp(\varphi^{\prime}(a+)z-Lz^{2}/2)\,\mathrm{d}z$
$=\ f(a)(b-a)^{2}\int_{0}^{1}u\exp\bigl{(}\varphi^{\prime}(a+)(b-a)u-cu^{2})\,\mathrm{d}u$
$=\ f(a)(b-a)^{2}N\bigl{(}\varphi^{\prime}(a+)(b-a)\bigr{)}\operatorname{\mathbb{E}}\exp(-cU^{2}),$
where $c:=L(b-a)^{2}/2$, and $U$ denotes a nonnegative random variable with density proportional to $1_{[u<1]}u\exp(\varphi^{\prime}(a+)(b-a)u)$. Similarly, $\mu(a,b)$ is not smaller than
$\int_{0}^{b-a}$ $z\exp(\psi_{4}(z))\,\mathrm{d}z\Big{/}\int_{0}^{b-a}\exp(\psi_{4}(z))\,\mathrm{d}z$
$\geq\ \int_{0}^{b-a}z\exp(\varphi^{\prime}(a+)z-Lz^{2}/2)\,\mathrm{d}z\Big{/}\int_{0}^{b-a}\exp(\varphi^{\prime}(a+)z)\,\mathrm{d}z$
$=\ (b-a)\int_{0}^{1}u\exp\bigl{(}\varphi^{\prime}(a+)(b-a)u-cu^{2})\,\mathrm{d}u\Big{/}\int_{0}^{1}\exp(\varphi^{\prime}(a+)(b-a)u)\,\mathrm{d}u$
$=\ (b-a)\nu\bigl{(}\varphi^{\prime}(a+)(b-a)\bigr{)}\operatorname{\mathbb{E}}\exp(-cU^{2}).$
But the random variable $U$ is stochastically smaller than a gamma random variable $Y$ with shape parameter $2$ and rate parameter $|\varphi^{\prime}(a+)|(b-a)$. Thus it follows from Jensen’s inequality and this comparison that
$\operatorname{\mathbb{E}}\exp(-cU^{2})\ \geq\ \exp(-c\operatorname{\mathbb{E}}(U^{2}))\ \geq\ \exp(-c\operatorname{\mathbb{E}}(Y^{2}))\ =\ \exp\bigl{(}-3L\varphi^{\prime}(a+)^{-2}\bigr{)},$
because $\operatorname{\mathbb{E}}(Y^{2})\ =\ 6\varphi^{\prime}(a+)^{-2}(b-a)^{-2}$. ∎
3.4 Exponential and maximal inequalities
In what follows, let $\hat{M}_{n},\hat{\mu}_{n},\hat{W}_{n}$ and $\hat{M}_{n}^{\mathrm{emp}},\hat{\mu}_{n}^{\mathrm{emp}},\hat{W}_{n}^{\mathrm{emp}}$ be defined as $M,\mu,W$ with $\hat{P}_{n}$ and $\hat{P}_{n}^{\mathrm{emp}}$, respectively, in place of $P$. The following basic result will be our key to bound the ratios $\hat{M}_{n}^{\mathrm{emp}}/M$ and $\hat{\mu}_{n}^{\mathrm{emp}}/\mu$.
Proposition 4.
Let $Q$ be a distribution on $0,\infty)$ with log-concave density and mean $\mu_{Q}>0$. Consider an i.i.d. sample $Y_{1},\ldots,Y_{m}$ drawn from $Q$. Then, for arbitrary $t<1$,
$\operatorname{\mathbb{E}}\exp\Bigl{(}t\sum_{i=1}^{m}\frac{Y_{i}}{\mu_{Q}}\Bigr{)}\ \leq\ (1-t)^{-m}.$
Proof of Proposition 4.
If $Q$ follows an exponential distribution, it is well known that the above inequality holds with equality.
Now suppose that $Q$ is an arbitrary distribution with log-density $\varphi$. Let $\tilde{\varphi}$ be the log-density of the exponential distribution $\tilde{Q}$ with mean $\mu_{Q}$, that is, $\tilde{\varphi}(x)=-x/\mu_{Q}-\log(\mu_{Q})$ for $x\geq 0$ and $\tilde{\varphi}(x)=$-$\infty$ for $x<0$. If $\varphi\not\equiv\tilde{\varphi}$, then there exist points $0<a<b<\infty$ such that $\varphi<\tilde{\varphi}$ on $\mathbb{R}\setminus[a,b]$ and $\varphi>\tilde{\varphi}$ on $(a,b)$. Indeed, by concavity of $\varphi-\tilde{\varphi}$ on $0,\infty)$, the set $\{\varphi>\tilde{\varphi}\}$ is an interval with infimum $a\geq 0$ and supremum $b\leq\infty$, and $\varphi<\tilde{\varphi}$ on $0,\infty)\setminus[a,b]$. Since $\int e^{\varphi(x)}\,\mathrm{d}x=\int e^{\tilde{\varphi}(x)}\,\mathrm{d}x=1$, we know that $a<b$ and $(a,b)\neq(0,\infty)$. If $a=0$, then the distribution function $F$ of $Q$ would be strictly larger than the distribution function $\tilde{F}$ of $\tilde{Q}$ on $(0,\infty)$, leading to the contradiction that $\mu_{Q}=\int_{0}^{\infty}(1-F(x))\,\mathrm{d}x<\mu_{Q}$. With the same argument one can argue that $b<\infty$. But now, for any convex function $\Psi$ on the real line,
$\int\Psi\,\mathrm{d}Q\ \leq\ \int\Psi\,\mathrm{d}\tilde{Q}$. (11)
Indeed, since $\int(c+dx)\,P(\mathrm{d}x)=\int(c+dx)\,\tilde{P}(\mathrm{d}x)$ for arbitrary $c,d\in\mathbb{R}$, the difference $\int\Psi\,dQ-\int\Psi\,\mathrm{d}\tilde{Q}$ equals
$\int\Psi(x)(e^{\varphi(x)}-e^{\tilde{\varphi}(x)})\,\mathrm{d}x\ =\ \int\Bigl{(}\Psi(x)-\frac{b-x}{b-a}\Psi(a)-\frac{x-a}{b-a}\Psi(b)\Bigr{)}(e^{\varphi(x)}-e^{\tilde{\varphi}(x)})\,\mathrm{d}x\ \leq\ 0,$
because the integrand of the latter integral is non-positive. Applying (11) to $\Psi(x)=\exp(tx)$ for arbitrary $t<1$, and using independence of $Y_{1},\ldots,Y_{m}$ concludes the proof. ∎
Our next key results are simultaneous inequalities for the univariate versions of $\hat{\mu}_{n}^{\mathrm{emp}}/\mu$ and $\hat{\mu}_{n}/\mu$.
Proposition 5.
(a)  For any $\tau>1$, the probability that
$\Bigl{|}\frac{\hat{\mu}_{n}^{\mathrm{emp}}(X_{(k)})}{\mu(X_{(k)})}-1\Bigr{|}\ <\ \sqrt{\frac{2\tau\log(n)}{n-k}}+\frac{\tau\log(n)}{n-k}\ \ \text{for}\ k=1,\ldots,n-1$
is at least $1-2n^{1-\tau}$.
(b)  For arbitrary constants $b_{n}$ such that $b_{n}\to b_{o}$ and $(1-F(b_{n}))/\rho_{n}\to\infty$,
$\max_{k<n:X_{(k)}\leq b_{n}}\,\Bigl{|}\frac{\hat{\mu}_{n}^{\mathrm{emp}}(X_{(k)})}{\mu(X_{(k)})}-1\Bigr{|}\ =\ O_{\mathrm{p}}\bigl{(}\sqrt{\rho_{n}/(1-F(b_{n})}\bigr{)},$
$\max_{x\in\hat{\mathcal{S}}_{n}:x\leq b_{n}}\,\Bigl{|}\frac{\hat{\mu}_{n}(x)}{\mu(x)}-1\Bigr{|}\ =\ O_{\mathrm{p}}\bigl{(}\sqrt{\rho_{n}/(1-F(b_{n})}\bigr{)}.$
Proof of Proposition 5.
As to part (a), suppose first that $P$ is the exponential distribution with mean $\mu(0)$. Then Chernov’s bound, applied to exponential distributions, shows that for $\varepsilon\geq 0$,
$\operatorname{\mathbb{P}}\biggl{(}\pm\Bigl{(}\frac{\hat{\mu}_{n}^{\mathrm{emp}}(0)}{\mu(0)}-1\Bigr{)}\geq\varepsilon\biggr{)}\ \leq\ \exp[-nH(\pm\varepsilon)],$ (12)
where $H(t):=t-\log(1+t)$ for t>1t>-1 and $H(t):=\infty$ for $t\leq-1$. Due to Proposition 4 the Chernov bound (12) holds true for arbitrary distributions $P$ with log-concave density such that $a_{o}\geq 0$. Coming back to the general case, note that for any $k\in\{1,\ldots,n-1\}$ and $a<b_{o}$, the conditional distribution of $(X_{(k+\ell)}-a)_{\ell=1}^{n-k}$, given that $X_{(k)}=a$, coincides with the distribution of $(Y_{(\ell)})_{\ell=1}^{n-k}$, where $Y_{(1)}\leq\cdots\leq Y_{(n-k)}$ are the order statistics of nkn-k independent random variables with density $f_{a}(y):=1_{[y\geq 0]}f(a+y)/(1-F(a))$. Since $\hat{\mu}_{n}^{\mathrm{emp}}(a)$ is the mean of $X_{(k+\ell)}-a$, $1\leq\ell\leq n-k$, we may apply the inequalities (12) to deduce that
$\operatorname{\mathbb{P}}\biggl{(}\pm\Bigl{(}\frac{\hat{\mu}_{n}^{\mathrm{emp}}(X_{(k)})}{\mu(X_{(k)})}-1\Bigr{)}\geq\varepsilon\biggr{)}\ \leq\ \exp[-(n-k)H(\pm\varepsilon)]$ (13)
for arbitrary $\varepsilon\geq 0$. Since $H(-\varepsilon)\geq H(\varepsilon)$ for all $\varepsilon\geq 0$, this implies that
$\operatorname{\mathbb{P}}\biggl{(}\Bigl{|}\frac{\hat{\mu}_{n}^{\mathrm{emp}}(X_{(k)})}{\mu(X_{(k)})}-1\Bigr{|}\geq\varepsilon\biggr{)}\ \leq\ 2\exp[-(n-k)H(\varepsilon)].$ (14)
Note that $H\bigl{(}\sqrt{2r}+r\bigr{)}\geq r$ for arbitrary $r\geq 0$, because $H\bigl{(}\sqrt{2r}+r\bigr{)}-r$ is equal to $\sqrt{2r}-\log\bigl{(}1+\sqrt{2r}+r\bigr{)}$, and $\exp\bigl{(}\sqrt{2r}\bigr{)}\geq 1+\sqrt{2r}+\sqrt{2r}^{2}/2=1+\sqrt{2r}+r$. Consequently,
$\operatorname{\mathbb{P}}$ $\biggl{(}\Bigl{|}\frac{\hat{\mu}_{n}^{\mathrm{emp}}(X_{(k)})}{\mu(X_{(k)})}-1\Bigr{|}\geq\sqrt{\frac{2\tau\log(n)}{n-k}}+\frac{\tau\log(n)}{n-k}\ \ \text{for some}\ k\in\{1,\ldots,n-1\}\biggr{)}$
$\leq\ \sum_{k=1}^{n-1}\operatorname{\mathbb{P}}\biggl{(}\Bigl{|}\frac{\hat{\mu}_{n}^{\mathrm{emp}}(X_{(k)})}{\mu(X_{(k)})}-1\Bigr{|}\geq\sqrt{\frac{2\tau\log(n)}{n-k}}+\frac{\tau\log(n)}{n-k}\biggr{)}$
$\leq\ 2\sum_{k=1}^{n-1}\exp(-\tau\log(n))\ <\ 2n^{1-\tau}.$
Concerning part (b), note that $\log(n)/(n-k)$ equals $\rho_{n}/(1-\hat{F}_{n}^{\mathrm{emp}}(X_{(k)})$, so combining part (a) with (6) shows that the maximum of $\bigl{|}\hat{\mu}_{n}^{\mathrm{emp}}(X_{(k)})/\mu(X_{(k)})-1\bigr{|}$ over all $k$ such that $X_{(k)}\leq b_{n}$ is of order $O_{\mathrm{p}}\bigl{(}\sqrt{\rho_{n}/(1-F(b_{n})}\bigr{)}$. Combining part (a) with (4), (6) and (7) shows that the maximum of $\bigl{|}\hat{\mu}_{n}(x)/\mu(x)-1\bigr{|}$ over all $x\in\hat{\mathcal{S}}_{n}$ such that $x\leq b_{n}$ is of order $O_{\mathrm{p}}\bigl{(}\sqrt{\rho_{n}/(1-F(b_{n})}\bigr{)}$ too. ∎
Finally, we need some inequalities for the ratios $\hat{M}_{n}^{\mathrm{emp}}(a,b)/M(a,b)$, $\hat{M}_{n}(a,b)/M(a,b)$ over a broad range of pairs $(a,b)$.
Proposition 6.
(a)  For any real number $a$ and $b\in(a,\infty]$ such that $P(a,b)>0$,
$\operatorname{\mathbb{P}}\biggl{(}\Bigl{|}\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b)}{M(a,b)}-1\Bigr{|}\geq 2\sqrt{\frac{\tau}{P(a,b)}}+\frac{\tau}{P(a,b)}\biggr{)}\ \leq\ 2e^{-n\tau}$
for all $\tau>0$.
(b)  For any sequence of numbers $\delta_{n}\in(0,1)$ such that $\delta_{n}\to 0$ and $\delta_{n}/\rho_{n}\to\infty$, let $\mathcal{A}_{n}:=\bigl{\{}(a,b):-\infty<a<b\leq\infty,P(a,b)\geq\delta_{n}\bigr{\}}$. Then
$\sup_{(a,b)\in\mathcal{A}_{n}}\,\Bigl{|}\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b)}{M(a,b)}-1\Bigr{|}\ \to_{\mathrm{p}}\ 0.$
Moreover,
$\sup_{(a,b)\in\mathcal{A}_{n}:b\in\hat{\mathcal{S}}_{n}}\,\Bigl{(}\frac{\hat{M}_{n}(a,b)}{M(a,b)}-1\Bigr{)}^{+}\ \to_{\mathrm{p}}\ 0.$
Proof of Proposition 6.
As to part (a), we first note that Proposition 4 yields the inequality
$\operatorname{\mathbb{E}}\exp\Bigl{(}t\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b)}{M(a,b)}\Bigr{)}\$$=\ \operatorname{\mathbb{E}}\exp\Bigl{(}\frac{t}{nP(a,b)}\sum_{i=1}^{n}1_{(a,b)}(X_{i})\frac{X_{i}-a}{\mu(a,b)}\Bigr{)}$
$=\ \operatorname{\mathbb{E}}\operatorname{\mathbb{E}}\Bigl{(}\exp\Bigl{(}\frac{t}{nP(a,b)}\sum_{i=1}^{n}1_{(a,b)}(X_{i})\frac{X_{i}-a}{\mu(a,b)}\Bigr{)}\,\Big{|}\,\hat{P}_{n}^{\mathrm{emp}}(a,b)\Bigr{)}$
$\leq\ \sum_{k=0}^{\infty}\binom{n}{k}P(a,b)^{k}(1-P(a,b))^{n-k}\Bigl{(}1-\frac{t}{nP(a,b)}\Bigr{)}^{-k}$
$=\ \biggl{(}1+\frac{P(a,b)t}{nP(a,b)-t}\biggr{)}^{n}$
$\leq\ \exp\Bigl{(}\frac{nP(a,b)t}{nP(a,b)-t}\Bigr{)}\ =\ \exp\Bigl{(}t+\frac{t^{2}}{nP(a,b)-t}\Bigr{)}$
for arbitrary $t<nP(a,b)$, so
$\operatorname{\mathbb{E}}\exp\Bigl{(}t\Bigl{(}\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b)}{M(a,b)}-1\Bigr{)}\Bigr{)}\ \leq\ \exp\Bigl{(}\frac{t^{2}}{nP(a,b)-t}\Bigr{)}.$
A standard application of Markov’s inequality shows that for $\eta\geq 0$,
$\operatorname{\mathbb{P}}\biggl{(}\Bigl{|}\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b)}{M(a,b)}-1\Bigr{|}\geq\eta\biggr{)}\ \leq\ 2\exp\Bigl{(}\frac{t^{2}}{nP(a,b)-t}-t\eta\Bigr{)}$
for all $t\in[0,nP(a,b))$. This bound is minimized for $t=nP(a,b)\bigl{(}1-1/\sqrt{1+\eta}\bigr{)}$, which leads to
$\operatorname{\mathbb{P}}\biggl{(}\Bigl{|}\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b)}{M(a,b)}-1\Bigr{|}\geq\eta\biggr{)}\ \leq\ 2\exp\Bigl{(}-nP(a,b)\bigl{(}\sqrt{1+\eta}-1\bigr{)}^{2}\Bigr{)}.$
Thus, for $\eta=2\sqrt{\tau/P(a,b)}+\tau/P(a,b)$ we obtain the asserted inequality.
Concerning part (b), let $\mathcal{D}_{n}=\bigl{\{}j/n:j\in\mathbb{Z}\cap[-n^{2},n^{2}]\bigr{\}}\cup\{\infty\}$. Then it follows from part (a) that for $D>4$,
$\operatorname{\mathbb{P}}\biggl{(}$ $\Bigl{|}\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b)}{M(a,b)}-1\Bigr{|}\geq 2\sqrt{\frac{D\rho_{n}}{P(a,b)}}+\frac{D\rho_{n}}{P(a,b)}\ \text{for some}\ a,b\in\mathcal{D}_{n},P(a,b)>0\biggr{)}$
$\leq\ (2n^{2}+1)2n^{2}n^{-D}\ \to\ 0.$
In particular,
$\Delta_{n,M}\ :=\ \max_{a,b\in\mathcal{D}_{n}:P(a,b)\geq\delta_{n}/2}\,\Bigl{|}\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b)}{M(a,b)}-1\Bigr{|}\ \to_{\mathrm{p}}\ 0.$
Note also that by Proposition 1,
$\Delta_{n,P}\ :=\ \max_{a,b\in\mathcal{D}_{n}:P(a,b)\geq\delta_{n}/2}\,\Bigl{|}\frac{\hat{P}_{n}^{\mathrm{emp}}(a,b)}{P(a,b)}-1\Bigr{|}\ \to_{\mathrm{p}}\ 0.$
For an arbitrary pair $(a,b)\in\mathcal{A}_{n}$, let $(a^{\prime},a^{\prime\prime}]$ and $(b^{\prime},b^{\prime\prime}]$ be minimal intervals with endpoints in $\{-\infty\}\cup\mathcal{D}_{n}$ containing $a$ and $b$, respectively. Note that $P(a^{\prime},a^{\prime\prime})$ and $P(b^{\prime},b^{\prime\prime})$ are not larger than
$\gamma_{n}\ :=\ \max\bigl{(}\{P(c,c+1/n):c\in\mathbb{R}\}\cup\{P(-\infty,-n),P(n,\infty)\}\bigr{)}\ =\ O(n^{-1}),$
because $P$ has a bounded density and subexponential tails. In particular, $P(a^{\prime\prime},b^{\prime})\geq\delta_{n}-2\gamma_{n}\geq\delta_{n}/2$ for sufficiently large $n$. On the one hand,
$\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b)}{M(a,b)}\ \leq\ \frac{\hat{M}_{n}^{\mathrm{emp}}(a,b^{\prime\prime})}{M(a,b^{\prime\prime})}\cdot\frac{M(a,b^{\prime\prime})}{M(a,b)}$
and
$\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b)}{M(a,b)}\ \geq\ \frac{\hat{M}_{n}^{\mathrm{emp}}(a,b^{\prime})}{M(a,b^{\prime})}\cdot\frac{M(a,b^{\prime})}{M(a,b)},$
and by Proposition 3 (b), the ratios $M(a,b^{\prime\prime})/M(a,b)$ and $M(a,b)/M(a,b^{\prime})$ are larger than one but not larger than $p_{n}+(1-p_{n})\log(1-p_{n})]^{-1}\to 1$, where $p_{n}:=(\delta_{n}-\gamma_{n})^{+}/\delta_{n}\to 1$. On the other hand, it follows from Fubini’s theorem that uniformly in $(a,b)\in\mathcal{A}_{n}$ and for sufficiently large $n$,
$\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b^{\prime\prime})}{M(a,b^{\prime\prime})}\$$=\ \frac{\int_{a}^{a^{\prime\prime}}\hat{P}_{n}^{\mathrm{emp}}(x,b^{\prime\prime})\,\mathrm{d}x+\hat{M}_{n}^{\mathrm{emp}}(a^{\prime\prime},b^{\prime\prime})}{\int_{a}^{a^{\prime\prime}}P(x,b^{\prime\prime})\,\mathrm{d}x+M(a^{\prime\prime},b^{\prime\prime})}$
$\leq\ \frac{(1+\Delta_{n,P})\int_{a}^{a^{\prime\prime}}P(x,b^{\prime\prime})\,\mathrm{d}x+(1+\Delta_{n,M})M(a^{\prime\prime},b^{\prime\prime})}{\int_{a}^{a^{\prime\prime}}P(x,b^{\prime\prime})\,\mathrm{d}x+M(a^{\prime\prime},b^{\prime\prime})}$
$\leq\ 1+\max\{\Delta_{n,P},\Delta_{n,M}\}$
and
$\frac{\hat{M}_{n}^{\mathrm{emp}}(a,b^{\prime})}{M(a,b^{\prime})}\$$=\ \frac{\int_{a}^{a^{\prime\prime}}\hat{P}_{n}^{\mathrm{emp}}(x,b^{\prime})\,\mathrm{d}x+\hat{M}_{n}^{\mathrm{emp}}(a^{\prime\prime},b^{\prime})}{\int_{a}^{a^{\prime\prime}}P(x,b^{\prime\prime})\,\mathrm{d}x+M(a^{\prime\prime},b^{\prime\prime})}$
$\geq\ \frac{(1-\Delta_{n,P})\int_{a}^{a^{\prime\prime}}P(x,b^{\prime})\,\mathrm{d}x+(1-\Delta_{n,M})M(a^{\prime\prime},b^{\prime})}{\int_{a}^{a^{\prime\prime}}P(x,b^{\prime})\,\mathrm{d}x+M(a^{\prime\prime},b^{\prime})}$
$\geq\ 1-\max\{\Delta_{n,P},\Delta_{n,M}\}.$
These considerations show that the supremum of $\bigl{|}\hat{M}_{n}^{\mathrm{emp}}(a,b)/M(a,b)-1\bigr{|}$ over all $(a,b)\in\mathcal{A}_{n}$ converges to 0 in probability.
Concerning the ratio $\hat{M}_{n}(a,b)/M(a,b)$, note first that for any probability distribution $Q$ with finite first moment and real numbers $a<b$,
$\int(x-a)^{+}\,Q(\mathrm{d}x)-\int(x-b)^{+}\,Q(\mathrm{d}x)\ =\ \int_{(a,b]}(x-a)Q(\mathrm{d}x)+(b-a)Q(b,\infty).$
Consequently, it follows from (4) and (5) that
$\lim_{b^{\prime}\to b+}\hat{M}_{n}^{\mathrm{emp}}(a,b^{\prime})\ \geq\ \hat{M}_{n}(a,b)\quad\text{if}\ b\in\hat{\mathcal{S}}_{n}.$
This inequality implies the assertion about the ratio $\hat{M}_{n}(a,b)/M(a,b)$. ∎
By symmetry the following results are an immediate consequence of Proposition 6.
Corollary 7.
For any sequence of numbers $\delta_{n}\in(0,1)$ such that $\delta_{n}\to 0$ and $\delta_{n}/\rho_{n}\to\infty$, let $\tilde{\mathcal{A}}_{n}:=\bigl{\{}(a,b):-\infty\leq a<b<\infty,P(a,b)\geq\delta_{n}\bigr{\}}$. Then
$\sup_{(a,b)\in\tilde{\mathcal{A}}_{n}}\,\Bigl{|}\frac{\hat{W}_{n}^{\mathrm{emp}}(a,b)}{W(a,b)}-1\Bigr{|}\ \to_{\mathrm{p}}\ 0.$
Moreover,
$\sup_{(a,b)\in\tilde{\mathcal{A}}_{n}:a\in\hat{\mathcal{S}}_{n}}\,\Bigl{(}\frac{\hat{W}_{n}(a,b)}{W(a,b)}-1\Bigr{)}^{+}\ \to_{\mathrm{p}}\ 0.$
4 Proofs of the main results
Proof of Theorem 1.
Since $f$ and $\hat{f}_{n}$ are zero on $\mathbb{R}\setminus(a_{o},b_{o})$, and because of (2), it suffices to show that for fixed points $a_{o}<a<b<b_{o}$,
$\sup_{x\in(a_{o},a]}\,\bigl{(}\hat{f}_{n}(x)-f(x)\bigr{)}\ \leq\ L(a)+o_{\mathrm{p}}(1)\quad\text{and}\quad\sup_{x\in[b,b_{o})}\,\bigl{(}\hat{f}_{n}(x)-f(x)\bigr{)}\ \leq\ R(b)+o_{\mathrm{p}}(1)$
with bounds $L(a),R(b)$ such that $L(a)\to 0$ as $a\downarrow a_{o}$ and $R(b)\to 0$ as $b\uparrow b_{o}$. For symmetry reasons we only consider the second claim. We fix an arbitrary $m\in(a_{o},b_{o})$ such that $\varphi^{\prime}(m+)<0$ in case of $\varphi^{\prime}(b_{o}-)<0$ and restrict our attention to $b\in(m,b_{o})$. If $b_{o}<\infty$ and $\varphi^{\prime}(b_{o}-)\geq 0$, then concavity of $\varphi$ and $\hat{\varphi}_{n}$ implies that
$\sup_{x\in[b,b_{o})}\,\bigl{(}\hat{f}_{n}(x)-f(x)\bigr{)}\$$\leq\ \sup_{x\in[b,b_{o})}\,\hat{f}_{n}(b)\exp\Bigl{(}\frac{\hat{\varphi}_{n}(b)-\hat{\varphi}_{n}(m)}{b-m}(x-b)\Bigr{)}-f(b)$
$\to_{\mathrm{p}}\ f(b)\bigl{(}f(b)/f(m)\bigr{)}^{(b_{o}-b)/(b-m)}-f(b)$
$\leq\ f(b_{o})\bigl{[}\bigl{(}f(b_{o})/f(m)\bigr{)}^{(b_{o}-b)/(b-m)}-1\bigr{]}\ =:\ R(b),$
and $R(b)\to 0$ as $b\uparrow b_{o}$. If $\varphi^{\prime}(b_{o}-)<0$, then
$\sup_{x\in[b,b_{o})}\,\bigl{(}\hat{f}_{n}(x)-f(x)\bigr{)}\$$\leq\ \sup_{x\in[b,b_{o})}\,\hat{f}_{n}(b)\exp\Bigl{(}\frac{\hat{\varphi}_{n}(b)-\hat{\varphi}_{n}(m)}{b-m}(x-b)\Bigr{)}-f(b_{o})$
$\to_{\mathrm{p}}\ f(b)-f(b_{o})\ =:\ R(b),$
because $\bigl{(}\hat{\varphi}_{n}(b)-\hat{\varphi}_{n}(m)\bigr{)}/(b-m)\to_{\mathrm{p}}\bigl{(}\varphi(b)-\varphi(m)\bigr{)}/(b-m)\leq\varphi^{\prime}(m+)<0$, and $R(b)\to 0$ as $b\uparrow b_{o}$. Here $f(\infty):=0$.
Let $(b_{n})_{n}$ be a sequence in $(a_{o},b_{o})$ with limit $b_{o}$. For arbitrary fixed $a_{o}<a<b<b_{o}$, it follows from concavity of $\hat{\varphi}_{n}$ that for sufficiently large $n$,
$\hat{\varphi}_{n}^{\prime}(b_{n}+)\ \leq\ \frac{\hat{\varphi}_{n}(b)-\hat{\varphi}_{n}(a)}{b-a}\ \to_{\mathrm{p}}\ \frac{\varphi(b)-\varphi(a)}{b-a}.$
Now the assertion follows from the fact that the right-hand side converges to $\varphi^{\prime}(b_{o}-)$ as $a,b\to b_{o}$. ∎
Proof of Theorem 2.
As to part (a), it follows from (2) and Theorem 1 that for any fixed $b\in(a,b_{o})$,
$\sup_{x\geq a}\,\bigl{|}\hat{f}_{n}(x)-f(x)\bigr{|}\$$\leq\ \sup_{x\in[a,b]}\,\bigl{|}\hat{f}_{n}(x)-f(x)\bigr{|}+\sup_{x\geq b}\,\bigl{|}\hat{f}_{n}(x)-f(x)\bigr{|}$
$\leq\ o_{\mathrm{p}}(1)+\sup_{x\geq b}\,\bigl{(}\hat{f}_{n}(x)-f(x)\bigr{)}^{+}+\sup_{x\geq b}\,f(x)$
$=\ o_{\mathrm{p}}(1)+\sup_{x\geq b}\,f(x),$
and $\sup_{x\geq b}f(x)\to f(b_{o})=0$ as $b\uparrow b_{o}$. Furthermore, since $\varphi^{\prime}(b_{o}-)=$-$\infty$, it follows from Theorem 1 that $\hat{\varphi}_{n}^{\prime}(b_{n}+)\to_{\mathrm{p}}\varphi^{\prime}(b_{o}-)$.
As to part (b), it follows from (1) and Theorem 1 that for any fixed $b\in(a,b_{o})$ and all $n$ with $b_{n}>b$,
$\sup_{x\in[a,b_{n}]}\,\bigl{|}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{|}\$$\leq\ \sup_{x\in[a,b]}\,\bigl{|}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{|}+\sup_{x\in[b,b_{n}]}\,\bigl{|}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{|}$
$\leq\ o_{\mathrm{p}}(1)+\sup_{x\in[b,b_{o}]}\,\bigl{(}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{)}^{+}+\sup_{x\in[b,b_{n}]}\,\bigl{(}\varphi(x)-\hat{\varphi}_{n}(x)\bigr{)}^{+}$
$=\ o_{\mathrm{p}}(1)+\sup_{x\in[b,b_{n}]}\,\bigl{(}\varphi(x)-\hat{\varphi}_{n}(x)\bigr{)}^{+},$
where we used the fact that $\delta(b):=\min_{x\in[b,b_{o}]}f(x)>0$, so $(\hat{\varphi}_{n}-\varphi)^{+}\leq(\hat{f}_{n}-f)^{+}/\delta(b)$ on $b,b_{o}]$. By concavity of $\hat{\varphi}_{n}$,
$\sup_{x\in[b,b_{n}]}\bigl{(}\varphi(x)-\hat{\varphi}_{n}(x)\bigr{)}^{+}$ $\leq\sup_{x\in[b,b_{n}]}\bigl{(}\varphi(x)-\varphi(b_{o})\bigr{)}^{+}+\sup_{x\in[b,b_{n}]}\bigl{(}\varphi(b_{o})-\hat{\varphi}_{n}(x)\bigr{)}^{+}$
$=\sup_{x\in[b,b_{o}]}\,\bigl{(}\varphi(x)-\varphi(b_{o})\bigr{)}^{+}+\max_{x\in\{b,b_{n}\}}\,\bigl{(}\varphi(b_{o})-\hat{\varphi}_{n}(x)\bigr{)}^{+}$
$\leq 2\sup_{x\in[b,b_{o}]}\bigl{|}\varphi(x)-\varphi(b_{o})\bigr{|}+\bigl{(}\varphi(b)-\hat{\varphi}_{n}(b)\bigr{)}^{+}+\bigl{(}\varphi(b_{o})-\hat{\varphi}_{n}(b_{n})\bigr{)}^{+}$
$=2\sup_{x\in[b,b_{o}]}\bigl{|}\varphi(x)-\varphi(b_{o})\bigr{|}+o_{\mathrm{p}}(1)+\bigl{(}\varphi(b_{o})-\hat{\varphi}_{n}(b_{n})\bigr{)}^{+}$
by (1). Since $\sup_{x\in[b,b_{o}]}\bigl{|}\varphi(x)-\varphi(b_{o})\bigr{|}\to 0$ as $b\uparrow b_{o}$, it suffices to show that
$\bigl{(}\varphi(b_{o})-\hat{\varphi}_{n}(b_{n})\bigr{)}^{+}\ \to_{\mathrm{p}}\ 0.$
To this end, we show that for any fixed $\varepsilon>0$, the inequality $\hat{\varphi}_{n}(b_{n})\leq\varphi(b_{o})-\varepsilon$ holds with asymptotic probability zero. Let $b(\varepsilon)\in(a_{o},b_{o})$ such that $|\varphi-\varphi(b_{o})|\leq\lambda\varepsilon/2$ on $b(\varepsilon),b_{o}]$ for some $\lambda\in(0,1)$ to be specified later. From (1) and Theorem 1 we may conclude that $\hat{\varphi}_{n}(b(\varepsilon))\geq\varphi(b_{o})-\lambda\varepsilon$ und $\hat{\varphi}_{n}\leq\varphi(b_{o})+\lambda\varepsilon$ on $b(\varepsilon),b_{o}]$ with asymptotic probability one. Thus it suffices to show that the event
$A_{n,\varepsilon}\ :=\ \bigl{[}\hat{\varphi}_{n}(b_{n})\leq\varphi(b_{o})-\varepsilon,\hat{\varphi}_{n}(b(\varepsilon))\geq\varphi(b_{o})-\lambda\varepsilon,\hat{\varphi}_{n}\leq\varphi(b_{o})+\lambda\varepsilon\ \text{on}\ [b(\varepsilon),b_{o}]\bigr{]}$
has asymptotic probability zero. From now on we assume that the event $A_{n,\varepsilon}$ occurs. Suppose that $n$ is sufficiently large such that $b_{n}>b(\varepsilon)$. Note that $\hat{\varphi}_{n}^{\prime}(b_{n}+)<0$, because $\hat{\varphi}_{n}(b(\varepsilon))>\hat{\varphi}_{n}(b_{n})$. Let $Y_{n}$ be the largest point in $\hat{\mathcal{S}}_{n}\cap(a_{o},b_{n}]$. Then, $\hat{\varphi}_{n}$ is affine on $Y_{n},b_{n}]$ and non-increasing on $Y_{n},b_{o}]$. Consequently, $\hat{f}_{n}$ is convex on $Y_{n},b_{n}]$ and non-increasing on $Y_{n},b_{o}]$.
Suppose first that $Y_{n}\geq b(\varepsilon)$. Then the properties of $\hat{f}_{n}$ on $Y_{n},b_{o}]$ imply that
$1-\hat{F}_{n}(Y_{n})\ =\ \hat{P}_{n}([Y_{n},b_{o}])\$$\leq\ (b_{n}-Y_{n})\frac{\hat{f}_{n}(Y_{n})+\hat{f}_{n}(b_{n})}{2}+(b_{o}-b_{n})\hat{f}_{n}(b_{n})$
$\leq\ (b_{o}-Y_{n})\frac{\hat{f}_{n}(Y_{n})+\hat{f}_{n}(b_{n})}{2}$
$\leq\ (b_{o}-Y_{n})f(b_{o})\frac{e^{\lambda\varepsilon}+e^{-\varepsilon}}{2},$
whereas
$1-F(Y_{n})\ =\ P([Y_{n},b_{o}])\ \geq\ (b_{o}-Y_{n})f(b_{o})e^{-\lambda\varepsilon/2}.$
Consequently, for sufficiently large $n$, the event $A_{n,\varepsilon}$ implies that
$\frac{1-\hat{F}_{n}(Y_{n})}{1-F(Y_{n})}\ \leq\ \frac{e^{3\lambda\varepsilon/2}+e^{\lambda\varepsilon/2-\varepsilon}}{2}\ =\ 1+(\lambda-1/2)\varepsilon+O(\varepsilon^{2})$
as $\varepsilon\downarrow 0$. Hence, if $\lambda<1/2$ and $\varepsilon>0$ is sufficiently small, it follows from (7) that the event $A_{n,\varepsilon}$ has asymptotic probability zero.
Suppose that $Y_{n}\leq b(\varepsilon)$. Then,
$\hat{P}_{n}([b(\varepsilon),b_{o}])\$$\leq\ (b_{n}-b(\varepsilon))\frac{\hat{f}_{n}(b(\varepsilon))+\hat{f}_{n}(b_{n})}{2}+(b_{o}-b_{n})\hat{f}_{n}(b_{n})$
$\leq\ (b_{o}-b(\varepsilon))f(b_{o})\frac{e^{\lambda\varepsilon}+e^{-\varepsilon}}{2},$
whereas
$P([Y_{n},b_{o}])\ \geq\ (b_{o}-b(\varepsilon))f(b_{o})e^{-\lambda\varepsilon/2}.$
Consequently, for sufficiently large $n$, the event $A_{n,\varepsilon}$ implies that
$(P-\hat{P}_{n})([b(\varepsilon),b_{o}])\$$\geq\ (b_{o}-b(\varepsilon))f(b_{o})\Bigl{(}e^{-\lambda\varepsilon/2}-\frac{e^{\lambda\varepsilon}+e^{-\varepsilon}}{2}\Bigr{)}$
$=\ (b_{o}-b(\varepsilon))f(b_{o})\bigl{(}(1/2-\lambda)\varepsilon+O(\varepsilon^{2})\bigr{)}$
as $\varepsilon\downarrow 0$. Hence, if $\lambda<1/2$ and $\varepsilon>0$ is sufficiently small, it follows from (3) that the event $A_{n,\varepsilon}$ has asymptotic probability zero.
As to part (c), because of Theorem 1, it suffices to show that for any fixed $\varepsilon>0$, the event $B_{n,\varepsilon}:=\bigl{[}\hat{\varphi}_{n}^{\prime}(b_{n}+)<\varphi^{\prime}(b_{o}-)-\varepsilon\bigr{]}$ has asymptotic probability zero. Let $Y_{n}$ be the largest point in $\hat{\mathcal{S}}_{n}$ such that $Y_{n}\leq b_{n}$. It follows from (1) that for any fixed $a\in(a_{o},b_{o})$, the event $B_{n,\varepsilon}\cap[Y_{n}\leq a]$ has asymptotic probability zero, because for any fixed $b\in(a,b_{o})$,
$\hat{\varphi}_{n}^{\prime}(a+)\ \geq\ \frac{\hat{\varphi}_{n}(b)-\hat{\varphi}_{n}(a)}{b-a}\ \to_{\mathrm{p}}\ \frac{\varphi(b)-\varphi(a)}{b-a}\ \geq\ \varphi^{\prime}(b_{o}-),$
whereas $\hat{\varphi}_{n}^{\prime}(Y_{n}+)=\hat{\varphi}_{n}^{\prime}(b_{n}+)$. Consequently, there exist numbers $a_{n,\varepsilon}\in(a_{o},b_{n})$ such that $a_{n,\varepsilon}\to b_{o}$ and $\operatorname{\mathbb{P}}(B_{n,\varepsilon}\cap[Y_{n}\leq a_{n,\varepsilon}])\to 0$. It remains to be shown that $B_{n,\varepsilon}\cap[Y_{n}>a_{n,\varepsilon}]$ has asymptotic probability zero. Assuming that the latter event occurs, note that by Proposition 3 (c),
$\mu(Y_{n})\$$\geq\ (b_{o}-Y_{n})\nu\bigl{(}\varphi^{\prime}(b_{o}-)(b_{o}-Y_{n})\bigr{)},$
$\hat{\mu}_{n}(Y_{n})\$$\leq\ (b_{o}-Y_{n})\nu\bigl{(}(\varphi^{\prime}(b_{o}-)-\varepsilon)(b_{o}-Y_{n})\bigr{)},$
and since the moduli of $\varphi^{\prime}(b_{o}-)(b_{o}-Y_{n})$ and $(\varphi^{\prime}(b_{o}-)-\varepsilon)(b_{o}-Y_{n})$ are not larger than $\bigl{(}|\varphi^{\prime}(b_{o}-)|+\varepsilon\bigr{)}(b_{o}-a_{n,\varepsilon})\to 0$, we may conclude that
$\frac{\hat{\mu}_{n}(Y_{n})}{\mu(Y_{n})}-1\$$\leq\ \frac{1/2+(\varphi^{\prime}(b_{o}-)-\varepsilon)(b_{o}-Y_{n})/12+O(1)(b_{o}-Y_{n})^{2}}{1/2+\varphi^{\prime}(b_{o}-)(b_{o}-Y_{n})/12+O(1)(b_{o}-Y_{n})^{2}}-1$
$=\ -\frac{\varepsilon(b_{o}-Y_{n})/6+O(1)(b_{o}-Y_{n})^{2}}{1+\varphi^{\prime}(b_{o}-)(b_{o}-Y_{n})/6+O(1)(b_{o}-Y_{n})^{2}}$
$=\ -(\varepsilon+o(1))(b_{o}-Y_{n})/6$
$\leq\ -(\varepsilon+o(1))(b_{o}-b_{n})/6$
$=\ -\bigl{(}\varepsilon/f(b_{o})+o(1)\bigr{)}(1-F(b_{n})),$
uniformly on $B_{n,\varepsilon}\cap[Y_{n}>a_{n,\varepsilon}]$. On the other hand, we know from Proposition 5 (b) that
$\frac{\hat{\mu}_{n}(Y_{n})}{\mu(Y_{n})}-1\ \geq\ O_{\mathrm{p}}\bigl{(}\sqrt{\rho_{n}/(1-F(b_{n}))}\bigr{)}\ =\ o_{\mathrm{p}}(1-F(b_{n})),$
because $\rho_{n}/(1-F(b_{n}))^{3}\to 0$ by assumption, whence $\operatorname{\mathbb{P}}(B_{n,\varepsilon}\cap[Y>a_{n,\varepsilon}])\to 0$. ∎
Proof of Theorem 3.
Concerning part (a), since $\{\hat{f}_{n}>0\}=[X_{(1)},X_{(n)}]$,
$\operatorname{\mathbb{P}}(\hat{f}_{n}(b_{n})=0)\ =\ \operatorname{\mathbb{P}}(X_{(1)}>b_{n})+\operatorname{\mathbb{P}}(X_{(n)}<b_{n})\ =\ (1-F(b_{n}))^{n}+F(b_{n})^{n}\ \to\ 0,$
because $F(b_{n})\to 1$ and $n(1-F(b_{n}))\to\infty$. If $\varphi^{\prime}(\infty-)=$-$\infty$, it follows already from Theorem 1 that $\hat{\varphi}_{n}^{\prime}(b_{n}+)\to_{\mathrm{p}}\varphi^{\prime}(\infty-)$. Otherwise, we know that $\hat{\varphi}_{n}^{\prime}(b_{n}+)\leq\varphi(\infty-)+o_{\mathrm{p}}(1)$, and it suffices to show that for any fixed $\varepsilon>0$, the inequality $\hat{\varphi}_{n}^{\prime}(b_{n}+)\leq\varphi^{\prime}(\infty-)-\varepsilon$ holds true with asymptotic probability zero. If $\hat{\varphi}_{n}^{\prime}(b_{n}+)\leq\varphi^{\prime}(\infty-)-\varepsilon$, then it follows from Proposition 3 (d) that $Y_{n}:=\max(\hat{\mathcal{S}}_{n}\cap(a_{o},b_{n}])$, satisfies
$\mu(Y_{n})\$$\geq\ -1/\varphi^{\prime}(\infty-),$
$\hat{\mu}_{n}(Y_{n})\$$\leq\ -1/\hat{\varphi}_{n}^{\prime}(Y_{n}+)\ =\ -1/\hat{\varphi}_{n}^{\prime}(b_{n}+)\ \leq\ -1/(\varphi^{\prime}(\infty-)-\varepsilon),$
whence
$\frac{\hat{\mu}_{n}(Y_{n})}{\mu(Y_{n})}\ \leq\ \frac{\varphi^{\prime}(\infty-)}{\varphi^{\prime}(\infty-)-\varepsilon}\ =\ \bigl{(}1+\varepsilon/|\varphi^{\prime}(\infty-)|\bigr{)}^{-1}.$
According to Proposition 5 (b), the latter inequality holds true with asymptotic probability zero.
To prove the claim in part (b), recall that for any compact interval $I\subset(a_{o},\infty)$,
$\Delta_{n}(I)\ :=\ \max_{x\in I}\,\bigl{|}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{|}\ \to_{\mathrm{p}}\ 0.$
Consequently, there exists a sequence $(a_{n})_{n}$ of numbers $a_{n}\in[a,b_{n}]$ converging to $\infty$ such that even $\Delta_{n}[a,a_{n}]\to_{\mathrm{p}}0$. Hence, it suffices to show that
$\max_{x\in[a_{n},b_{n}]}\,\frac{\bigl{|}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{|}}{1+|\varphi(x)|}\ \to_{\mathrm{p}}\ 0.$
We may assume without loss of generality that $\varphi(a_{n})\leq 0$ and $\varphi^{\prime}(a_{n}+)<0$. Applying part (a) to $(b_{n})_{n}$ and to $(a_{n})_{n}$ in place of $(b_{n})_{n}$ implies that
$\Gamma_{n}\ :=\ \max_{t\in[a_{n},b_{n}]}\,\bigl{|}\hat{\varphi}_{n}^{\prime}(t+)/\varphi^{\prime}(t+)-1\bigr{|}\ \to_{\mathrm{p}}\ 0.$
Thus, uniformly in $x\in[a_{n},b_{n}]$,
$\bigl{|}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{|}\$$\leq\ \bigl{|}\hat{\varphi}_{n}(a_{n})-\varphi(a_{n})\bigr{|}+\int_{a_{n}}^{x}\bigl{|}\hat{\varphi}_{n}^{\prime}(t+)-\varphi^{\prime}(t+)\bigr{|}\,\mathrm{d}t$
$\leq\ o_{\mathrm{p}}(1)+\Gamma_{n}\int_{a_{n}}^{x}|\varphi^{\prime}(t+)|\,\mathrm{d}t$
$=\ o_{\mathrm{p}}(1)+\Gamma_{n}\bigl{(}\varphi(a_{n})-\varphi(x)\bigr{)}$
$\leq\ o_{\mathrm{p}}(1)+\Gamma_{n}|\varphi(x)|.$
Now we prove the claims in part (c). If $a<b<\infty$ and and $0<\delta<a-a_{o}$, then for $x\in[a,b]$,
$\hat{\varphi}_{n}^{\prime}(x+)-\varphi^{\prime}(x)\$$\leq\ \frac{\hat{\varphi}_{n}(x)-\hat{\varphi}_{n}(x-\delta)}{\delta}-\varphi^{\prime}(x)$
$\leq\ 2\Delta_{n}[a-\delta,b]/\delta+\frac{1}{\delta}\int_{x-\delta}^{x}\bigl{(}\varphi^{\prime}(s)-\varphi^{\prime}(x)\bigr{)}\,\mathrm{d}s$
$\leq\ 2\Delta_{n}[a-\delta,b]/\delta+L\delta/2,$
where $L$ denotes the Lipschitz constant of $\varphi^{\prime}$ on $(a_{*},\infty)$. Analogously,
$\hat{\varphi}_{n}^{\prime}(x+)-\varphi^{\prime}(x)\ \geq\ -2\Delta_{n}[a,b+\delta]/\delta-L\delta/2,$
Hence, $\max_{x\in[a,b]}\bigl{|}\hat{\varphi}_{n}^{\prime}(x+)-\varphi^{\prime}(x)\bigr{|}\to_{\mathrm{p}}0$. Consequently, there exist sequences $(a_{n})_{n}$ and $(\varepsilon_{n})_{n}$ of numbers $a_{n}\in[a,b_{n}]$ and $\varepsilon_{n}>0$ such that $a_{n}\to\infty$, $\varepsilon_{n}\to 0$, and with asymptotic probability one,
$\max_{x\in[a,a_{n}]}\,\bigl{|}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{|}\ <\ \varepsilon_{n},\quad\sup_{x\in[a,a_{n}]}\,\bigl{|}\hat{\varphi}_{n}^{\prime}(x+)-\varphi^{\prime}(x)\bigr{|}\ <\ \varepsilon_{n}.$ (15)
Consequently, it suffices to verify the claims of part (c) with $a_{n},b_{n}]$ in place of $a,b_{n}]$.
At first we show that
$\Gamma_{n}\ :=\ \sup_{x\in[a_{n},b_{n}]}\,\Bigl{(}\frac{\hat{\varphi}_{n}^{\prime}(x+)}{\varphi^{\prime}(x)}-1\Bigr{)}^{+}\ \to_{\mathrm{p}}\ 0.$ (16)
It follows from Proposition 3 (e,d,a,d) that
$\sup_{x\in[a_{n},b_{n}]}\,\frac{\hat{\varphi}_{n}^{\prime}(x+)}{\varphi^{\prime}(x)}\ =\ (1+o(1))\sup_{x\in[a_{n},b_{n}]}\,\mu(x)|\hat{\varphi}_{n}^{\prime}(x+)|\ \leq\ (1+o(1))\max_{y\in\hat{\mathcal{S}}_{n}:y\leq b_{n}}\,\frac{\mu(y)}{\hat{\mu}_{n}(y)},$
because for any $x\in[a_{n},b_{n}]$, the point $y:=\max(\hat{\mathcal{S}}_{n}\cap(a_{o},x])$ satisfies $y\leq x$ and $\hat{\varphi}_{n}^{\prime}(y+)=\hat{\varphi}_{n}^{\prime}(x+)\leq\hat{\varphi}_{n}^{\prime}(a_{n}+)<0$ with asymptotic probability one, whence
$\mu(x)|\hat{\varphi}_{n}^{\prime}(x+)|\ \leq\ \mu(y)|\hat{\varphi}_{n}^{\prime}(x+)|\ =\ \mu(y)|\hat{\varphi}_{n}^{\prime}(y+)|\ \leq\ \mu(y)/\hat{\mu}_{n}(y).$
Now (16) follows from the fact that the maximum of $\mu/\hat{\mu}_{n}$ over $\hat{\mathcal{S}}_{n}\cap(a_{o},b_{n}]$ equals $1+o_{\mathrm{p}}(1)$, see Proposition 5 (b).
Secondly, with (16) at hand, the claim about $(\hat{\varphi}_{n}-\varphi)^{+}/(1+|\varphi|)$ can be verified easily. Uniformly in $x\in[a_{n},b_{n}]$,
$\varphi(x)-\hat{\varphi}_{n}(x)\$$=\ \varphi(a_{n})-\hat{\varphi}_{n}(a_{n})+\int_{a_{n}}^{x}\bigl{(}\varphi^{\prime}(t)-\hat{\varphi}_{n}^{\prime}(t+)\bigr{)}\,\mathrm{d}t$
$\leq\ \varepsilon_{n}+\Gamma_{n}\int_{a_{n}}^{x}|\varphi^{\prime}(t)|\,\mathrm{d}t$
$=\ \varepsilon_{n}+\Gamma_{n}\bigl{(}\varphi(a_{n})-\varphi(x)\bigr{)}$
$\leq\ o_{\mathrm{p}}(1)\bigl{(}1+|\varphi(x)|\bigr{)}.$
Finally, we derive an upper bound for $(\hat{\varphi}_{n}-\varphi)^{+}$ on $a_{n},b_{n}]$. To this end we augment the boundary $a_{n}$ by a number $a_{n}^{\prime}\in[a,a_{n}]$ such that
$\varphi^{\prime}(a_{n}^{\prime})\ =\ \min\{\varphi^{\prime}(a),\varphi^{\prime}(a_{n})+1\}.$
Since $\varphi^{\prime}(a_{n}^{\prime})\to$-$\infty$, the sequence $(a_{n}^{\prime})_{n}$ converges to $\infty$ too. Moreover, assuming that (15) holds true, as soon as $\varphi^{\prime}(a_{n}^{\prime})=\varphi^{\prime}(a_{n})+1$ and $\varepsilon_{n}<1/2$,
$\hat{\varphi}_{n}^{\prime}(a_{n}^{\prime}+)-\hat{\varphi}_{n}^{\prime}(a_{n}+)\ \geq\ 1-2\varepsilon_{n}\ >\ 0,$
so
$\operatorname{\mathbb{P}}\bigl{(}\hat{\mathcal{S}}_{n}\cap[a_{n}^{\prime},a_{n}]\neq\emptyset)\ \to\ 1.$ (17)
For any $x\geq a$ let
$y(x)\ :=\ x+|\varphi^{\prime}(x)|^{-1}.$
This defines an increasing function $y:[a,\infty)\to(a,\infty)$ such that
$\frac{P(x,y(x))}{P(x,\infty)}\ \to\ 1-e^{-1}\quad\text{as}\ x\to\infty.$ (18)
Indeed,
$\frac{P(x,y(x))}{P(x,\infty)}\$$=\ \int_{0}^{y(x)-x}\exp\bigl{(}\varphi(x+s)-\log\bar{F}(x)\bigr{)}\,\mathrm{d}s$
$\geq\ |\varphi^{\prime}(x)|\int_{0}^{y(x)-x}\exp(-|\varphi^{\prime}(x)|s)\,\mathrm{d}s$
= $=\ 1-e^{-1},$
because the distribution with density $s\mapsto\exp\bigl{(}\varphi(x+s)-\log\bar{F}(x)\bigr{)}$ on $(0,\infty)$ is stochastically smaller than the exponential distribution with rate $|\varphi^{\prime}(x)|$. On the other hand, since $|\varphi^{\prime}(x)|s\geq\varphi(x+s)-\varphi(x)\geq-|\varphi^{\prime}(x)|s-Ls^{2}/2$ for $s\geq 0$,
$\frac{P(x,y(x))}{P(x,\infty)}\$$\leq\ \int_{0}^{y(x)-x}\exp\bigl{(}-|\varphi^{\prime}(x)|s\bigr{)}\,\mathrm{d}s\Big{/}\int_{0}^{\infty}\exp\bigl{(}-|\varphi^{\prime}(x)|s-Ls^{2}/2\bigr{)}\,\mathrm{d}s$
$=\ (1-e^{-1})\Big{/}\int_{0}^{\infty}e^{-y}\exp(-L\varphi^{\prime}(x)^{-2}y^{2}/2\bigr{)}\,\mathrm{d}y$
$\leq\ \exp(L\varphi^{\prime}(x)^{-2})(1-e^{-1})$
by Jensen’s inequality.
Now we define $b_{n}^{\prime}:=y(b_{n})$ and $b_{n}^{\prime\prime}:=y(b_{n}^{\prime})$. Then it follows from (18) that
$\frac{1-F(b_{n}^{\prime\prime})}{\rho_{n}}\ =\ (e^{-2}+o(1))\frac{1-F(b_{n})}{\rho_{n}}\ \to\ \infty$
Hence, (16) remains valid with $a_{n}^{\prime},b_{n}^{\prime\prime}]$ in place of $a_{n},b_{n}]$. Furthermore, (18) implies that
$\inf_{x\in[a_{n}^{\prime},b_{n}^{\prime}]}\frac{P(x,y(x))}{\rho_{n}}\ =\ (1-e^{-1}+o(1))\frac{1-F(b_{n}^{\prime})}{\rho_{n}}\ \to\ \infty.$
Consequently, by Corollary 7,
$\max_{x\in[a_{n}^{\prime},b_{n}^{\prime}]\cap\hat{\mathcal{S}}_{n}}\,\frac{\hat{W}_{n}(x,y(x))}{W(x,y(x))}\ \leq\ 1+o_{\mathrm{p}}(1).$
On the other hand, Proposition 3 (c) implies that for $x\in[a_{n}^{\prime},b_{n}^{\prime}]$,
$\hat{W}_{n}(x,y(x))\$$\geq\ \hat{f}_{n}(x)\varphi^{\prime}(x)^{-2}V\bigl{(}-\hat{\varphi}_{n}^{\prime}(y(x)+)/\varphi(x)\bigr{)},$
$W(x,y(x))\$$\leq\ f(x)\varphi^{\prime}(x)^{-2}V(-1),$
because $\varphi^{\prime}<0$ on $a,\infty)$, and it follows from (16) with $a_{n}^{\prime},b_{n}^{\prime\prime}]$ in place of $a_{n},b_{n}]$, Lipschitz-continuity of $\varphi^{\prime}$ on $a,\infty)$ with constant $L$ and Proposition 2 that
$\inf_{x\in[a_{n}^{\prime},b_{n}^{\prime}]}\,\frac{V\bigl{(}-\hat{\varphi}_{n}^{\prime}(y(x)+)/\varphi(x)\bigr{)}}{V(-1)}\$$\geq\ \inf_{x\in[a_{n}^{\prime},b_{n}^{\prime}]}\,\frac{V\bigl{(}-(1+o_{\mathrm{p}}(1))\varphi^{\prime}(y(x))/\varphi^{\prime}(x)\bigr{)}}{V(-1)}$
$\geq\ \frac{V\bigl{(}-(1+o_{\mathrm{p}}(1))(1+L\varphi^{\prime}(a_{n}^{\prime})^{-2})\bigr{)}}{V(-1)}$
= $=\ 1+o_{\mathrm{p}}(1),$
because $\varphi^{\prime}(y(x))/\varphi^{\prime}(x)\geq-1-L\varphi^{\prime}(x)^{-2}\geq-1-L\varphi^{\prime}(a_{n}^{\prime})^{-2}\to-1$. Consequently,
$\max_{x\in[a_{n}^{\prime},b_{n}^{\prime}]\cap\hat{\mathcal{S}}_{n}}\,\frac{\hat{f}_{n}(x)}{f(x)}\ \leq\ 1+o_{\mathrm{p}}(1),$
which is equivalent to
$\max_{x\in[a_{n}^{\prime},b_{n}^{\prime}]\cap\hat{\mathcal{S}}_{n}}\,\bigl{(}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{)}^{+}\ =\ o_{\mathrm{p}}(1).$
Since $\hat{\varphi}_{n}-\varphi$ is convex between consecutive knots in $\hat{\mathcal{S}}_{n}$, and since $\hat{\mathcal{S}}_{n}\cap[a_{n}^{\prime},a_{n}]\neq\emptyset$ with asymptotic probability one by (17), this implies that
$\max_{x\in[a_{n},\hat{b}_{n}]}\,\bigl{(}\hat{\varphi}_{n}(x)-\varphi(x)\bigr{)}^{+}\ =\ o_{\mathrm{p}}(1),$
where $\hat{b}_{n}:=\max(\hat{\mathcal{S}}_{n}\cap(-\infty,b_{n}^{\prime}])$. This proves already our claim in case of $\hat{b}_{n}\geq b_{n}$. So it remains to show that
$\frac{\hat{f}_{n}(b_{n})}{f(b_{n})}\ \leq\ 1+o_{\mathrm{p}}(1)\quad\text{if}\ \hat{\mathcal{S}}_{n}\cap[b_{n},b_{n}^{\prime}]=\emptyset.$
With $\check{b}_{n}:=\min(\hat{\mathcal{S}}_{n}\cap[b_{n}^{\prime},\infty))$, the inequality $\hat{b}_{n}\leq b_{n}$ implies that $\hat{\varphi}_{n}$ is linear on $b_{n},\check{b}_{n}]$, and $P(b_{n},\check{b}_{n})/\rho_{n}\geq P(b_{n},b_{n}^{\prime})/\rho_{n}\geq(1-e^{-1}+o(1))(1-F(b_{n}))/\rho_{n}\to\infty$. Hence Proposition 6 (b) implies that
$\frac{\hat{M}_{n}(b_{n},\check{b}_{n})}{M(b_{n},\check{b}_{n})}\ \leq\ 1+o_{\mathrm{p}}(1).$
On the other hand, Proposition 3 (c,e) implies that
$\hat{M}_{n}(b_{n},\check{b}_{n})\$$\geq\ \hat{f}_{n}(b_{n})(\check{b}_{n}-b_{n})^{2}N\bigl{(}\hat{\varphi}_{n}^{\prime}(b_{n}+)(\check{b}_{n}-b_{n})\bigr{)},$
$M(b_{n},\check{b}_{n})\$$\leq\ f(b_{n})(\check{b}_{n}-b_{n})^{2}N\bigl{(}\varphi^{\prime}(b_{n})(\check{b}_{n}-b_{n})\bigr{)},$
and it follows from (16), Lipschitz-continuity of $\varphi^{\prime}$ on $a,\infty)$ with constant $L$ and Proposition 2 that
$\frac{N\bigl{(}\hat{\varphi}_{n}^{\prime}(b_{n}+)(\check{b}_{n}-b_{n})\bigr{)}}{N\bigl{(}\varphi^{\prime}(b_{n})(\check{b}_{n}-b_{n})\bigr{)}}\ \geq\ \frac{N\bigl{(}(1+o_{\mathrm{p}}(1))\varphi^{\prime}(b_{n})(\check{b}_{n}-b_{n})\bigr{)}}{N\bigl{(}\varphi^{\prime}(b_{n})(\check{b}_{n}-b_{n})\bigr{)}}\ =\ 1+o_{\mathrm{p}}(1).$
The latter conclusion follows from the fact that for arbitrary sequences $(t_{n})_{n}$ in $(-\infty,0]$, $N\bigl{(}(1+o(1))t_{n}\bigr{)}/N(t_{n})=1+o(1)$. For a bounded sequence $(t_{n})_{n}$ this follows from continuity of $N$. If $t_{n}\to$-$\infty$, this follows from the expansion $N\bigl{(}(1+o(1))t_{n}\bigr{)}=(1+o(1))t_{n}^{-2}$. ∎
Acknowledgement.
Part of this work was supported by the Swiss National Science Foundation.